### v1

# Exploring the evolution of red and blue galaxies in different cosmic web environments using IllustrisTNG simulation

Abstract
We analyze the evolution of red and blue galaxies in different cosmic web environments from redshift $z=3$ to $z=0$ using the IllustrisTNG simulation. We use Otsu’s method to classify the red or blue galaxies at each redshift and determine their geometric environments from the eigenvalues of the deformation tensor. Our analysis shows that initially, blue galaxies are more common in clusters followed by filaments, sheets and voids. However, this trend reverses at lower redshifts, with red fractions rising earlier in denser environments. At $z<1$, most massive galaxies ($\log(\frac{M_{*}}{M_{\odot}})>10.5$) are quenched across all environments. In contrast, low-mass galaxies ($\log(\frac{M_{*}}{M_{\odot}})<10.5$) are more influenced by their environment, with clusters hosting the highest red galaxy fractions at low redshifts. We observe a slower mass growth for low-mass galaxies in clusters at $z<1$. Filaments show relative red fractions (RRF) comparable to clusters at low masses, but host nearly $60\%$ of low-mass blue galaxies, representing a diverse galaxy population. It implies that less intense environmental quenching in filaments allows galaxies to experience a broader range of evolutionary stages. Despite being the densest environment, clusters display the highest relative blue fraction (RBF) for high-mass galaxies, likely due to interactions or mergers that can temporarily rejuvenate star formation in some of them. The $(u-r)$ colour distribution transitions from unimodal to bimodal by redshift $z=2$ across all environments. At $z<1$, clusters exhibit the highest median colour, with stellar mass being the primary driver of colour evolution in massive galaxies. The suppression of star formation rate (SFR) and specific SFR (sSFR) is also most pronounced in clusters during this period. Our study suggests that stellar mass governs quenching in high-mass galaxies, while a complex interplay of mass and environment shapes the evolution of low-mass galaxies.
1 Introduction
Galaxy colour bimodality [1, 2, 3, 4, 5] has been recognized for more than two decades. Observations from large galaxy surveys reveal that galaxies cluster into two distinct groups in colour-magnitude diagrams: a red, passive population and a blue, star-forming population. The red galaxies are typically older and have lower star formation rates. Their red colour indicates that they have largely exhausted their supply of gas and are composed of older, cooler stars. In contrast, the blue galaxies are actively forming new stars, which are hotter. These galaxies are generally younger and have a higher rate of star formation.
The implications of colour bimodality for galaxy formation and evolution are significant, as it provides insights into the life cycle of galaxies. Galaxies follow different evolutionary pathways during their evolution. The observed dichotomy suggests that galaxies undergo a transformation from blue to red over time. This transformation may result from external mechanisms such as interactions, mergers, and various environmental effects. Alternatively, internal processes, including gas depletion, stellar evolution, and feedback mechanisms, may also play crucial roles in this transformation. Understanding how different internal processes and external influences shape the observed bimodal colour distribution is essential for gaining a clear picture of galaxy evolution.
Blue galaxies dominate the cosmic landscape at earlier times due to the rising gas fractions at higher redshifts. Star formation in these galaxies is fueled by abundant gas reservoirs. Intense star formation activity depletes the available gas, leading to a decline in the star formation rate with time. As star formation gradually ceases with natural aging, the galaxy eventually becomes red, dominated by older and cooler stars. In addition to natural aging, various internal physical mechanisms, such as morphological quenching [6], mass quenching [7, 8, 9, 10], angular momentum quenching [11], and bar quenching [12], can also halt star formation in galaxies. Further, the expulsion or removal of gas from a galaxy provides another effective route for quenching star formation. Gas loss caused by feedback from supernovae, active galactic nuclei, shock-driven winds [14, 15, 16] and ram pressure stripping [13] can significantly suppress star formation in galaxies.
Besides the internal processes, environment play a crucial role in transforming blue galaxies into red galaxies. A large body of literature [17, 18, 4, 19, 20, 21, 23, 24, 22, 26, 27, 28, 29] shows that red and blue galaxies tend to inhabit high-density and low-density regions, respectively. Various environment-driven mechanisms, such as mergers [30], harassment [31, 32], strangulation [13, 33], starvation [34, 35, 36], and satellite quenching [37], can inhibit star formation and alter galaxy structures. Extensive research through simulations [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52] and observations [53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68] demonstrates that interactions between galaxies, which produce tidal torques, can trigger starbursts and change both the colour and morphology of galaxies.
The environment of a galaxy is often characterized by its local density, but this alone does not describe the neighbourhood of a galaxy. Galaxies are part of a complex cosmic web composed of interconnected filaments, clusters, sheets, and voids [69, 70, 71, 72, 73, 74, 75, 21, 76, 77], which collectively organizes and shapes the distribution of matter and galaxies on a grand scale. Studies with N-body simulations reveal a dynamic flow of matter within the cosmic web [78, 79, 80, 81], moving from voids to walls, walls to filaments, and finally into clusters.
Filaments are the largest known coherent structures in the universe [82, 83, 84]. Hydrodynamical simulations reveal [85, 86] that more than $40\%-50\%$ of baryonic matter is found in filaments as the Warm-Hot Intergalactic Medium (WHIM). This diffuse medium acts as a gas reservoir that can eventually fall into galaxies, contributing to their growth and affecting their star formation rates. The presence of WHIM can also influence gas accretion efficiency in galaxies. Galaxies in different regions of the cosmic web experience varying levels of gas accretion. For instance, galaxies near the centers of filaments and sheets receive a steady supply of cold gas, which fuels star formation and increases their mass [87, 28, 88, 68, 52, 89]. Clusters, the densest regions of the universe, typically form at the intersections of filaments and are characterized by frequent interactions with neighbouring galaxies and extreme environmental conditions, leading to relatively rapid transformations [13, 93, 94]. In contrast, galaxies in lower-density regions like sheets and voids generally follow quieter evolutionary paths with subdued star formation [95, 96]. Thus, different cosmic web environments play a crucial role in shaping the diverse evolutionary trajectories of galaxies across the universe.
Observations reveal that star formation activity in galaxies peaked around a redshift of $z\sim 2-3$ [97, 98], an era often dubbed the “cosmic noon”. The cosmic star formation rate has seen a dramatic decline from $z=1$ to $z=0$ [99]. The number of massive red galaxies with fixed stellar masses has steadily increased since $z\sim 1$ [100, 101]. These trends suggest significant changes in galaxy properties in recent times, which could be crucial in explaining the observed bimodality in galaxy distributions. It is essential to understand how the cosmic web influences the colour bimodality and its evolution.
The IllustrisTNG simulation [102, 103, 104, 105, 106, 107] models the formation and evolution of galaxies within the universe. It employs sophisticated numerical techniques to simulate the complex interactions of dark matter, gas, and stars across a vast volume of the universe. It incorporates detailed physics, including gas dynamics, star formation, supernova feedback, and black hole growth, to provide insights into galaxy formation and evolution. This simulation is ideal for studying the roles of different cosmic web environments on galaxy evolution. Several works [108, 109, 102, 110, 111, 112, 113] have explored quenching in galaxies using various hydrodynamical simulations. In the present work, we will utilize data from the IllustrisTNG simulation to investigate how different cosmic web environments influence the transformation of galaxy colour since a redshift of $z\sim 3$.
We will categorize the galaxies into red and blue populations based on their colour and stellar mass, following a recently introduced classification scheme [114] based on Otsu’s method [115]. We will investigate the evolution of the red and blue fractions in various cosmic web environments. Analyzing these fractions would help us to assess the efficiency of quenching processes within each environment. By tracking how the absolute numbers or percentages of red and blue galaxies change over time, we can understand how quenching impacts galaxy populations in different cosmic web environments. Additionally, we will examine the evolution of the relative red and blue fractions, which reveals how the proportion of red or blue galaxies in a given environment compares to those in all other cosmic environments. This comparison will reveal how the relative abundance of red and blue galaxies in each environment influences the development of colour bimodality. We will also explore how these fractions vary with the stellar mass of galaxies in different cosmic web environments. Studying how the stellar mass dependence of these fractions changes in different environments would help us understand the roles of both environment and stellar mass in shaping the colour bimodality. It is also important to understand how different galaxy properties evolve in different cosmic web environments. To explore this, we will examine the evolution of median colour, median stellar mass, median star formation rate (SFR), and median specific star formation rate (sSFR) of galaxies with redshift across various cosmic web environments.
The paper is organized into the following sections. In Section 2, we provide a detailed overview of the data used in this work. Section 3 explains the methodology employed for the analysis. The findings are discussed in Section 4, and Section 5 presents the conclusions drawn from our analysis.
2 Data
IllustrisTNG [^1] https://www.tng-project.org/ [106, 107, 102, 104, 105, 103] is a suite of cosmological gravo-magnetohydrodynamical simulations built on the moving-mesh code AREPO [116, 117]. It serves as an enhanced version of the original Illustris project [118, 119, 120]. The TNG suite consists of three cosmological volumes with side lengths of approximately 50, 100, and 300 Mpc, named TNG50, TNG100, and TNG300, respectively. Each volume offers multiple resolution levels, such as TNG100-1, TNG100-2, TNG100-3, and similarly for TNG300, whereas TNG50 has four distinct resolution outputs. All simulations begin at a redshift of $z=127$, using cosmological parameters from the Planck 2015 results [121]: $\Omega_{\Lambda}=0.6911$, $\Omega_{m}=0.3089$, $\Omega_{b}=0.0486$, $\sigma_{8}=0.8159$, $n_{s}=0.9667$, and $h=0.6774$. For our analysis, we utilize the TNG300-1 run, which offers the largest volume and highest resolution in the TNG series. The specifics of TNG300-1 are detailed in the  Table 1, and our study covers redshifts from $z=3$ to the present ($z=0$).
Simulation Name Volume $(\rm Mpc^{3})$ $N_{GAS}$ $N_{DM}$ $m_{baryon}\,(M_{\odot})$ $m_{DM}\,(M_{\odot})$
TNG300-1 $(302.6)^{3}$ $2500^{3}$ $2500^{3}$ $11\times 10^{6}$ $59\times 10^{6}$
Table 1: This table describes the specifications of the TNG300-1 simulation. Different columns in the table represent the volume of the box (comoving), initial number of gas cells ($N_{GAS}$), number of dark matter particles ($N_{DM}$), the target baryon mass ($m_{baryon}$) and dark matter particle mass ($m_{DM}$).
At each redshift, we extract various galaxy properties from the publicly available group catalog, which provides details on both halos and subhalos. The FoF algorithm is used to identify halos, while the subhalos (or galaxies) are identified using the SUBFIND algorithm [122, 123]. In this work, we select only galaxies with a non-zero SubhaloFlag, as a value of zero indicates a subhalo of non-cosmological origin. Additionally, we impose a stellar mass criterion, choosing galaxies in the range $129<\log_{10}(\frac{M_{\star}}{M_{\odot}})<$ 12. Stellar mass is defined as the total mass of star particles within twice the stellar half-mass radius $(r_{stars},1/2)$ [105]. For $(u-r)$ colour, we use the supplementary catalog provided by  [102]. The star formation rates are obtained from the SubhaloSFRinRad field in the group catalog. In IllustrisTNG, star formation is modeled following  [124]. For our analysis, we use the instantaneous star formation rate, which is the sum of the star formation rates of all star-forming gas cells within twice the stellar half-mass radius of the galaxy.
Figure 1: This figure shows the distribution of galaxies in the $(u-r)$ colour-stellar mass plane across each redshift, with the dividing line (determined by Otsu’s method) distinguishing the red and blue populations.
3 Method
The primary objective of this study is to classify galaxies in the IllustrisTNG simulation into red and blue populations according to their stellar mass and redshift, and to investigate how cosmic web environments and stellar mass influence their evolution.
3.1 Identifying red and blue Galaxies with Otsu’s method
Otsu’s thresholding technique [115], originally designed for image segmentation, separates data into two groups by minimizing the intra-class variance (variance within each group) and maximizing the inter-class variance (variance between groups). This method iterates through all possible thresholds, selecting the one that provides the optimal separation between the two classes. This technique is particularly useful when the data distribution is bimodal, as is the case with the $(u-r)$ colour distribution of galaxies, where we observe two peaks corresponding to the blue cloud and the red sequence. Recently, [114] proposed a method for classifying galaxies into two groups: the “blue cloud” (younger, star-forming galaxies) and the “red sequence” (older, passive galaxies) based on the Otsu’s technique.
The primary steps in this method are as follows.
  1. 1.
    Histogram Calculation: We first calculate the histogram of the $(u-r)$ colour for all galaxies in our sample, using a specific number of bins $M$. The histogram is then normalized by the total number of galaxies $N=\sum_{i=1}^{M}n_{i}$ where $n_{i}$ corresponds to the number of galaxies in the $i^{th}$ colour bin. This ensures that the sum of the probabilities for all bins equals to one i.e $\sum_{i=1}^{M}p_{i}=1$ where $p_{i}=\frac{n_{i}}{N}$.
  2. 2.
    Class Probabilities: For a given threshold, the galaxies are divided into two groups: blue cloud (BC) and red sequence (RS). The probability of each class occurring is calculated based on the proportion of galaxies in each group relative to the total sample. If the threshold is at the k-th bin, the blue cloud is represented by the first $k$ bins, while the red sequence is represented by the bins from $k+1$ to the end. The class probabilities are given by,
    $P_{BC}=\sum_{i=1}^{k}p_{i}=w(k)$ (3.1)
    and
    $P_{RS}=\sum_{i=k+1}^{M}p_{i}=1-w(k).$ (3.2)
  3. 3.
    Class means and Class variances: We calculate the mean colour and variance for both groups for each threshold value.
    The means of the blue cloud and the red sequence are calculated as,
    $\mu_{BC}=\frac{\sum_{i=1}^{k}x_{i}p_{i}}{P_{BC}}=\frac{\mu_{k}}{w(k)}$ (3.3)
    and
    $\mu_{RS}=\frac{\sum_{i=k+1}^{M}x_{i}p_{i}}{P_{RS}}=\frac{\mu_{T}-\mu_{k}}{1-w(k)}$ (3.4)
    Here, $x_{i}$ represents the $(u-r)$ colour corresponding to the $i^{th}$ bin. The mean up to the $k^{th}$ bin is given by $\mu_{k}=\sum_{i=1}^{k}x_{i}p_{i}$, while $\mu_{T}=\sum_{i=1}^{M}x_{i}p_{i}$ denotes the mean of the entire distribution. It is important to note that $P_{BC}+P_{RS}=1$ and that $\mu_{T}=P_{BC}\,\mu_{BC}+P_{RS}\,\mu_{RS}$ for every threshold.
    Similarly, the class variances are estimated as,
    $\sigma_{BC}^{2}=\frac{\sum_{i=1}^{k}(x_{i}-\mu_{BC})^{2}p_{i}}{P_{BC}}$ (3.5)
    and
    $\sigma_{RS}^{2}=\frac{\sum_{i=k+1}^{M}(x_{i}-\mu_{RS})^{2}p_{i}}{P_{RS}}$ (3.6)
  4. 4.
    Intra-class variance and Inter-class variance: The intra-class (within-class) variance ($\sigma_{wc}^{2}$) and the inter-class (between-class) variance ($\sigma_{bc}^{2}$) are calculated as,
    $\sigma_{wc}^{2}=P_{BC}\,\sigma_{BC}^{2}+P_{RS}\,\sigma_{RS}^{2}$ (3.7)
    and
    $\sigma_{bc}^{2}=P_{BC}\,P_{RS}\,(\mu_{BC}-\mu_{RS})^{2}$ (3.8)
    The total variance $\sigma_{T}^{2}$ is the sum of the two variances,
    $\sigma_{T}^{2}=\sigma_{wc}^{2}+\sigma_{bc}^{2}$ (3.9)
    It is important to note that both $\sigma_{wc}^{2}$ and $\sigma_{bc}^{2}$ are dependent on the selected threshold, while $\sigma_{T}^{2}$ remains independent of it.
  5. 5.
    Threshold Selection: The optimal threshold is the one that minimizes the intra-class variance $\sigma_{wc}^{2}$, which is equivalent to maximizing the inter-class variance $\sigma_{bc}^{2}$. In practice, this threshold can be found by iterating through the possible values and selecting the one that satisfies these conditions. This yields a threshold that is not affected by binning choices, making it a reliable approach for classifying red and blue galaxies [114]. This method offers a robust and parameter-free way to classify galaxies into the blue cloud and red sequence based on their $(u-r)$ colour and stellar mass.
Redshift Number of galaxies
All $9<\log_{10}\left(\frac{M_{\star}}{M_{\odot}}\right)\leq 10.5$ $1210.5<\log_{10}\left(\frac{M_{\star}}{M_{\odot}}\right)<$ 12
Total Red Blue Total Red Blue Total Red Blue
00 $253140$ $130751$ $122389$ $225075$ $105417$ $119658$ $28065$ $25334$ $2731$
$0.1$ $253212$ $120682$ $132530$ $225135$ $96243$ $128892$ $28077$ $24439$ $3638$
$0.2$ $252546$ $112362$ $140184$ $224612$ $89094$ $135518$ $27934$ $23268$ $4666$
$0.3$ $251329$ $105559$ $145770$ $223692$ $83375$ $140317$ $27637$ $22184$ $5453$
$0.4$ $249659$ $96073$ $153586$ $222588$ $75211$ $147377$ $27071$ $20862$ $6209$
$0.5$ $247760$ $90513$ $157247$ $221341$ $71415$ $149926$ $26419$ $19098$ $7321$
$0.7$ $242859$ $81067$ $161792$ $218375$ $64601$ $153774$ $24484$ $16466$ $8018$
$1$ $231933$ $73719$ $158214$ $210682$ $60488$ $150194$ $21251$ $13231$ $8020$
$1.5$ $199307$ $69796$ $129511$ $184161$ $61093$ $123068$ $15146$ $8703$ $6443$
$2$ $157394$ $66878$ $90516$ $147377$ $61740$ $85637$ $10017$ $5138$ $4879$
$3$ $84438$ $46230$ $38208$ $80992$ $44559$ $36433$ $3446$ $1671$ $1775$
Table 2: This table shows the total number of galaxies with nonzero SubhaloFlag and stellar masses within $129<\log_{10}(\frac{M_{\star}}{M_{\odot}})<$ 12 in IllustrisTNG simulation for each snapshot from redshift 00 to $3$. It also lists the number of red and blue galaxies identified with the Otsu’s method at each redshift in different mass ranges.
There are distinct relationships between colour and stellar mass or absolute magnitude. Therefore, a single colour threshold cannot be justified for galaxies of varying masses and luminosities. We divide the entire sample into several independent stellar mass bins and apply this technique to each of these bins separately. This provides us a dividing line between the two populations in the colour-stellar mass plane at each redshift (Figure 1). We also tabulate the number of red and blue galaxies in two different mass ranges between redshift $3$ to 00 in Table 2.
Figure 2: This figure displays the distribution of galaxies in various cosmic web environments at $z=0$, from a 20 Mpc thick slice of the IllustrisTNG simulation.
3.2 Identifying different morphological environments of the cosmic web
We classify galaxies into different morphological environments within the cosmic web using a Hessian-based method [125, 126]. This method relies on the eigenvalues and eigenvectors of the deformation tensor to identify voids, sheets, filaments, and clusters.
The deformation tensor $T_{ij}$ is derived from the Hessian matrix of the gravitational potential field $\Phi$, defined as:
$T_{ij}=\frac{\partial^{2}\Phi}{\partial x_{i}\partial x_{j}}$ (3.10)
where $x_{i}$ and $x_{j}$ are the spatial coordinates. The gravitational potential $\Phi$ is computed by solving the Poisson equation
$\nabla^{2}\Phi\equiv\delta$ (3.11)
Here, $\delta=\frac{\rho-\bar{\rho}}{\bar{\rho}}$ is the density contrast, with $\rho$ representing the local density and $\bar{\rho}$ the average density. To calculate the potential, we apply the Cloud-In-Cell (CIC) scheme to construct a discrete density contrast field on a grid. This field is smoothed using an isotropic Gaussian filter with a width of 4 Mpc.
Next, we compute the Fourier transform of the gravitational potential:
$\hat{\Phi}=\hat{G}\hat{\rho}$ (3.12)
where $\hat{G}$ is the Fourier transform of the Green’s function of the Laplacian operator, and $\hat{\rho}$ is the density in Fourier space. By transforming the potential back into real space, we calculate the tidal tensor using numerical differentiation.
Based on the signs of the three eigenvalues $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ (where $\lambda_{1}>\lambda_{2}>\lambda_{3}$), galaxies are classified into different cosmic web environments as follows:
  1. 1.
    Void: $\lambda_{1},\lambda_{2},\lambda_{3}<0$
  2. 2.
    Sheet: $\lambda_{1}>0$, $\lambda_{2},\lambda_{3}<0$
  3. 3.
    Filament: $\lambda_{1},\lambda_{2}>0$, $\lambda_{3}<0$
  4. 4.
    Cluster: $\lambda_{1},\lambda_{2},\lambda_{3}>0$
Each type represents a distinct geometric environment in the cosmic web, with voids being underdense and clusters representing the densest regions.
4 Results
4.1 Evolution of red and blue fractions in different cosmic web environments
In this subsection, we examine the evolution of the fraction of red and blue galaxies in various environments within the cosmic web. The red fraction (RF) and blue fraction (BF) in any cosmic web environment are respectively defined as $RF=\frac{n_{R}}{n_{R}+n_{B}}$ and $BF=\frac{n_{B}}{n_{R}+n_{B}}$. Here, $n_{R}$ and $n_{B}$ are the number of red and blue galaxies in the same environment. The top left and bottom left panels of Figure 3 illustrate these fractions across different cosmic web environments. Although the blue fraction can be obtained by simply subtracting the red fraction from one, they are separately shown here for the sake of completeness.
At redshift $z=3$, the blue fraction is highest in clusters, followed by filaments, sheets, and voids. Specifically, the blue fraction is $\sim 48\%$ in clusters and around $40\%$ in voids. Although the differences in blue fraction across environments at $z=3$ are not substantial, the trend suggests that overdense regions were more conducive to star formation in earlier times. This may arise due to the presence of larger gas reservoirs in these regions.
As redshift decreases, the blue fraction increases in all cosmic web environments. However, this trend reverses at lower redshifts, with the blue fraction declining and the red fraction rising. The blue fraction decreases and the red fraction increases at specific redshifts, with inflection points occurring at $z=1.5$ for clusters, $z=0.7$ for filaments and sheets, and $z=0.5$ for voids. At $z\leq 1.5$, red fractions dominate in clusters, followed by filaments, sheets, and voids. By $z=0$, the red fraction becomes $75\%$ in clusters, $45\%$ in filaments, $34\%$ in sheets, and $30\%$ in voids.
These results indicate an environment-dependent trend in galaxy colour transformation. Physical mechanisms that suppress star formation are more dominant in clusters and become effective earlier. This may be due to gradual gas depletion from ongoing star formation and faster gas exhaustion from interaction-induced starbursts in high-density regions. Suppression of star formation is less effective in filaments and least effective in sheets and voids, as evidenced by a high blue fraction ($\sim 68\%$) in these environments. This is also supported by a relatively smaller reduction in the blue fraction in sheets and voids from their respective peak values ($\sim 78\%$ at $z=0.5$).
[…]
Figure 5: Same as Figure 4 but for blue fraction.
We now divide our galaxy sample at each redshift into two distinct mass bins: $9<\log(\frac{M_{*}}{M_{\odot}})\leq 10.5$ and $1210.5<\log(\frac{M_{*}}{M_{\odot}})<$ 12 to examine the evolution of red and blue fractions in lower and higher mass galaxies separately. Results for these two mass bins are presented in the middle and right panels of Figure 3. The middle panels display the evolution of red and blue fractions for the lower mass bin, which includes approximately $\sim 90\%$ galaxies of the entire sample (Table 2). The results for the lower mass bin are quite similar to those for the entire sample for this reason.
The results for the higher mass bin, containing about $\sim 10\%$ of the galaxies, are shown in the right panels of Figure 3. Even for the galaxies in the higher mass bin at $z=3$, we observe a higher blue fraction and a lower red fraction in clusters compared to filaments and sheets. However, the red fraction steadily increases across all environments as redshift decreases, indicating that star formation is gradually suppressed in massive galaxies regardless of their environment. Notably, the blue fraction in voids increases between $z=2$ and $z=1$, suggesting that massive galaxies in voids may follow different evolutionary trajectories compared to those in other cosmic web environments.
4.2 Stellar mass dependence of the red and blue fractions in different cosmic web environments and their evolution
We also examine the red and blue fractions as functions of stellar mass within each geometric environment. Figure 4 and Figure 5 show these results across different redshifts. The bottom right panels of these figures reveal that at $z=3$, clusters generally have a higher blue fraction and a lower red fraction across nearly all masses. The blue fraction in filaments is lower than in clusters but higher than in sheets and voids throughout the entire mass range. However, the larger error bars for voids and sheets limit the statistical significance of these differences. Between $z=3-1.5$, we observe that in each geometric environment, the blue fraction increases significantly in lower mass galaxies ($\log(\frac{M_{*}}{M_{\odot}}\leq 10.5$). The blue fraction in lower mass galaxies do not exhibit a strong dependence on the geometric environment during this period. A clear environment-dependent trend emerges at $z<1$. We find that in clusters, the red fraction of lower mass galaxies increases from approximately $\sim 40\%$ to $\sim 70\%$ between redshifts 1 and 0. For higher mass galaxies, the red fraction rises from about $\sim 80\%$ to $\sim 95\%$ in all environments over the same period. Further, $\sim 40\%$ of galaxies in filaments are red, compared to around $\sim 30\%$ in sheets and voids. This indicates that the quenching of lower mass galaxies is strongly influenced by their environment at $z<1$, while quenching in more massive galaxies is less affected by their embedding geometric environments during the same period.
4.3 Evolution of relative red and blue fractions in different cosmic web environments
The red and blue fractions provide the proportion of two populations in any particular environment. We can use the relative red fraction (RRF) and the relative blue fraction (RBF) to study the relative abundance of red and blue galaxies in various cosmic web environments. The RRF and RBF are respectively defined as $RRF=\frac{(n_{R})_{i}}{\sum_{i=1}^{4}(n_{R})_{i}}$ and $RBF=\frac{(n_{B})_{i}}{\sum_{i=1}^{4}(n_{B})_{i}}$, where $(n_{R})_{i}$ and $(n_{B})_{i}$ are the numbers of red and blue galaxies in $i^{th}$-type environment. Here $i$ runs from $1$ to $4$ covering 4-types of cosmic web environment. We examine the evolution of the RRF and RBF across different cosmic web environments, as shown in the top and bottom left panels of Figure 6. The RRF and RBF represent the proportions of red and blue galaxies, respectively, within a specific environment compared to other cosmic environments. The top and bottom left panels of Figure 6 reveal that in filaments and sheets, the RRF decreases and the RBF increases with decreasing redshift. Conversely, clusters show an opposite trend where the RRF increases and the RBF decreases as redshift decreases. Throughout the entire redshift range, filaments consistently have the highest RBF. At lower redshifts ($z<1$), clusters exhibit the highest RRF compared to other environments.
The middle and right panels of Figure 6 present the RRF and RBF for lower and higher mass bins, respectively. The trends observed in the lower mass bin mirror those of the entire sample, given that most galaxies belong to the lower mass category. In the higher mass bin, the behavior of the RRF and RBF is less intuitive. Specifically, both RRF and RBF decrease in clusters with decreasing redshift, while they increase in filaments and sheets. For higher mass galaxies, clusters have higher RRF and RBF than filaments at earlier times ($z>1$), but this trend reverses at $z\sim 1$, with filaments surpassing clusters in both RRF and RBF. This counter-intuitive result can be explained by examining the relative proportions of galaxies in different cosmic web environments.
Figure 7 illustrates the evolution of the relative fractions of galaxies in filaments, sheets, clusters, and voids for the entire sample and the two distinct mass bins. The left panel of Figure 7 shows that filaments generally have the highest galaxy abundance, followed by clusters, sheets, and voids across the entire redshift range. The proportion of galaxies in filaments and sheets increases with decreasing redshift, while the proportion in clusters decreases. These trends are also evident in the lower mass bin, as depicted in the middle panel of Figure 7. In the high mass bin, the proportion of galaxies in filaments and clusters evolves differently, as shown in the right panel of Figure 7. There is a crossover around $z\sim 1$, after which filaments have a higher proportion of galaxies compared to clusters. Prior to $z=1$, clusters dominate in terms of galaxy proportion.
[…]
The anomalous results in the right panels of Figure 6 are explained by the trends observed in the right panel of Figure 7 and Figure 3. For high mass galaxies, RF increases and BF decreases in all types of cosmic web environments, with comparable values across different environments (Figure 3). Consequently, the magnitudes of the RRF and RBF are primarily determined by the proportions of galaxies in different environments. The trends observed in the right panel of Figure 7 are reflected in the top and bottom right panels of Figure 6, with crossovers in RRF and RBF occurring at different redshifts due to the varying red and blue fractions in filaments and clusters.
It may be noted that the red fraction (RB) and blue fraction (BF) are complementary to each other, meaning that their sum always equals 1. However, this relationship does not extend to the relative red fraction (RRF) and relative blue fraction (RBF). In other words, knowing the RRF does not allow us to directly infer the RBF.
4.4 Stellar mass dependence of the relative red and blue fractions in different cosmic web environments and their evolution
We now examine the RRF and RBF across different cosmic web environments as functions of stellar mass and their evolution from redshift $3$ to 00. The various panels in Figure 8 display the RRF as a function of stellar mass in different cosmic web environments at different redshifts. The results indicate that in both filaments and sheets, the RRF is lower at higher masses and higher at lower masses. The RRF in filaments and sheets at lower masses decreases with decreasing redshift. Conversely, clusters exhibit an opposite trend, where the RRF is higher at higher masses and lower at lower masses. In clusters, the RRF at higher masses remains relatively stable with redshift, while at lower masses, it gradually increases as redshift decreases. At $z\sim 1$, the RRF at lower masses becomes comparable between filaments and clusters. However, below this redshift, clusters tend to dominate at lower masses, while filaments become more prominent at intermediate masses. At low redshifts, the RRF in clusters exceeds that of all other environments at both low and high masses. This suggests that low-mass red galaxies in clusters may represent quenched satellites, while high-mass red galaxies may be massive centrals. Interestingly, the RRF in clusters and filaments shows two distinct crossovers at intermediate masses. Between these crossovers ($1110.5<\log(\frac{M_{*}}{M_{\odot}})<$ 11), the RRF in filaments surpasses that of all other environments. It may be worthwhile to mention here that the right panels of Figure 6 indicate that, at lower redshifts, both RRF and RBF dominate in filaments for masses $\log(\frac{M_{*}}{M_{\odot}})>10.5$. However, a closer examination reveals that this dominance is confined to the mass range $1110.5<\log(\frac{M_{*}}{M_{\odot}})<$ 11. There are only a small number of red and blue galaxies with masses $\log(\frac{M_{*}}{M_{\odot}})>11$, and these primarily reside in clusters.
[…]
Figure 9: Same as Figure 8 but for RBF.
We present the RBF as a function of stellar mass across different cosmic web environments and redshifts in the various panels of Figure 9. Our analysis reveals that in both filaments and sheets, the RBF is lower at higher masses and higher at lower masses. Specifically, the RBF in filaments and sheets at lower masses increases as redshift decreases. Between redshift $z=1-0$, filaments exhibit a significantly higher RBF ($\sim 0.55$) compared to other environments for galaxies having $\log(\frac{M_{*}}{M_{\odot}})<11$. This indicates that filaments host a diverse population of galaxies, encompassing both passive and star-forming galaxies. The less aggressive environmental quenching in filaments allows for a wider range of evolutionary stages. Interestingly, despite being the densest environment, clusters show the highest RBF for high-mass galaxies ($\log(\frac{M_{*}}{M_{\odot}}>11$) even at lower redshifts. This suggests that clusters are highly dynamic environments where interactions or mergers might temporarily reignite star formation activity in some high mass galaxies.
Figure 10: This figure shows the PDF of $(u-r)$ colour in different cosmic web environments at different redshifts.
[…]
4.5 Evolution of colour across different cosmic web environments
We present the distributions of $(u-r)$ colour across various cosmic web environments at different redshifts, ranging from $z=3$ to $z=0$, as shown in different panels of Figure 10. At $z=3$, the $(u-r)$ colour distribution in each environment is unimodal, with a peak around $(u-r)\sim 0.7$. As redshift decreases, this peak shifts towards higher $(u-r)$ values, indicating the aging of stellar populations. By $z=2$, the colour distribution begins to show bimodality. The bimodality becomes most pronounced in clusters, followed by filaments, sheets, and voids. In clusters, the peak corresponding to the red sequence grows rapidly, eventually surpassing the peak of the blue cloud in this environment at $z=0.5$. The emergence of distinct bimodality across all environments suggests that the colour transformation can occur in any cosmic web environment.
In the top left panel of Figure 11, we show the median $(u-r)$ colour of galaxies in various cosmic web environments as a function of redshift. Since $z\sim 2$, the median colour in clusters begins to diverge from that in other environments and increases rapidly after $z>1$. Meanwhile, the median colour in filaments starts to differ from that in sheets and voids after redshift 1. At $z<1$, clusters exhibit the highest median colour, followed by filaments, sheets, and voids.
The top right and bottom panels of Figure 11 illustrate the evolution of the median colour in each environment for lower and higher mass bins. The lower mass bin shows trends similar to the combined sample due to the large number of galaxies in this bin. The bottom panel reveals that the median colour for more massive galaxies ($1210.5<\log(\frac{M_{*}}{M_{\odot}})<$ 12) increases steadily at $z<1$, with little variation across different environments. This suggests that the colour evolution of the more massive galaxies is primarily driven by mass and is less dependent on their environment. This reaffirms our findings in subsection 4.1 and subsection 4.2.
Figure 12: Same as Figure 10 but for stellar mass.
Figure 13: Same as Figure 11 but for stellar mass.
4.6 Evolution of stellar mass across diverse cosmic web environments
We present the stellar mass distributions of galaxies across different environments from $z=3$ to $z=0$ in Figure 12. At higher masses, galaxies are predominantly found in clusters, followed by filaments, sheets, and voids. Conversely, at lower masses, the distribution is dominated by galaxies in voids, with fewer in sheets, filaments, and clusters. This trend indicates that massive galaxies are more common in clusters and filaments, while less massive galaxies are more prevalent in sheets and voids. At $z<2$, we observe the emergence of a bump in the stellar mass distribution around $\log(\frac{M_{*}}{M_{\odot}}\sim 10.5$) in each environment. The bump becomes more pronounced in all environments after $z=1$ suggesting an increase in the number of intermediate-mass galaxies during this period.
In the top left panel of Figure 13, we show the evolution of the median stellar mass of galaxies across different cosmic web environments. It shows that as redshift decreases, the median mass of galaxies increases in all environments. At each redshift, the median mass is highest in clusters, followed by filaments, sheets, and voids. The top right and bottom panels of Figure 13 show the results for lower and higher mass bins, respectively. In the top right panel, we observe that the median mass of galaxies in filaments and sheets increases steadily between $z=1$ and $z=0$. However, at $z<1$, the growth rate of the median mass for lower mass galaxies in clusters slows down significantly compared to those in filaments and sheets. The bottom panel shows that the median stellar mass for high mass galaxies remains stable between $z=3$ and $z=1.5$, followed by a gradual increase up to the present. This suggests that these galaxies undergo a period of minimal mass growth during the earlier stages of cosmic evolution, likely indicating that major mergers are less frequent at higher redshifts. The steady increase in median stellar mass at $z<1.5$ points to more significant mass accumulation in later stages, possibly due to mergers and accretion.
It may be interesting to note that while the median colour of massive galaxies is quite similar across different environments (as shown in Figure 11), their median mass differs significantly.
Figure 14: Same as Figure 10 but for SFR.
Figure 15: Same as Figure 11 but for SFR.
Figure 16: Same as Figure 10 but for sSFR.
Figure 17: Same as Figure 11 but for sSFR.
4.7 Evolution of SFR across diverse cosmic web environments
The SFR distributions of galaxies in various cosmic web environments are shown at different redshifts in the panels of Figure 14. The peak location of the SFR distribution shifts from $\log(SFR)=0.75$ to $\log(SFR)=-0.5$ over the redshift range $z=3$ to $z=0$. While the peak locations differ slightly across cosmic web environments between redshifts $z=3$ and $z=1$, they converge to nearly the same $\log(SFR)$ value at $z<1$. The peak amplitude is highest in voids, followed by sheets, filaments, and clusters.
We show the evolution of the median SFR in different environments in the three panels of Figure 15. The top left panel shows that the median SFR is highest in clusters, followed by filaments, sheets, and voids until $z=1.5$. The median SFR in clusters becomes smaller than the other geometric environments at $z<1$. The results for the lower mass bin, shown in the top right panel of Figure 15, exhibit a similar trend to that observed in the entire sample. The bottom panel of Figure 15 indicates that the median SFR for massive galaxies ($\log(\frac{M_{*}}{M_{\odot}})>10.5$) drops to zero at $z\sim 1.5$ across all environments, suggesting that star formation in high-mass galaxies is strongly suppressed in the low-redshift universe.
4.8 Evolution of sSFR across diverse cosmic web environments
We show the distribution of specific star formation rate (sSFR) across cosmic web environments at various redshifts in the panels of Figure 16. While the star formation rate (SFR) measures the degree of star formation in a galaxy, sSFR normalizes this rate by its stellar mass. This normalization is crucial for comparing star formation activity across galaxies of different sizes, offering insights into the efficiency of star formation relative to galaxy mass. The panels in Figure 16 reveal trends similar to those seen in Figure 14. Specifically, the peak of the sSFR distribution in each environment shifts towards more negative values as redshift decreases. This gradual shift points to both internal and external mechanisms leading to the aging and transformation of galaxies in each type of environment. The peak amplitude is highest in voids, followed by sheets, filaments, and clusters. This indicates that, for a given stellar mass, star formation is strongest in less dense environments and becomes progressively weaker as the environment grows denser.
The median sSFR across different environments is shown in Figure 17. The top left panel indicates that, unlike the median SFR, there are no statistically significant differences in the median sSFR across environments in the redshift range $31.5<z<$ 3. As redshift decreases, sSFR decreases in all environments. At $z<1$, clusters exhibit the lowest median sSFR, with filaments showing somewhat higher sSFR compared to clusters between $z=1$ and $z=0$. Sheets and voids have the highest and similar median sSFR in this redshift range. The top right panel presents results for the lower mass bin, which mirrors the trends observed in the entire sample for the reasons discussed earlier. The bottom panel of Figure 17 displays results for the higher mass bin. This panel shows similar trends but with larger differences in median sSFR across environments. It is interesting to note that the median sSFR for massive galaxies in voids tends to increase between $z=2$ and $z=1$, though this trend is difficult to confirm due to large errorbars for the galaxies in voids.
5 Discussions and Conclusions
We study the evolution of red and blue galaxies in different cosmic web environments using IllustrisTNG simulation. Our key findings are summarized as follows.
(i) We track the evolution of the red and blue fractions across different cosmic web environments from redshift $z=3$ to $z=0$. Initially, blue galaxies are more prevalent in clusters compared to other geometric environments. This trend reverses at lower redshifts. The red fraction starts to rise after specific redshifts for each environment, with clusters experiencing earlier suppression of star formation compared to filaments, sheets, and voids. The massive galaxies show a consistent increase in red fractions over time in all environments.
(ii) We analyze the red and blue fractions as a function of stellar mass across different environments at various redshifts. At $z=3$, clusters have a higher blue fraction and lower red fraction compared to other environments at all masses, with filaments showing an intermediate blue fraction. Over time, the red fraction increases in higher mass galaxies and decreases in lower mass galaxies at all environments until $z=1.5$. A pronounced environment-dependent trend emerges at $z<1$ during which all environments show a significant increase in the red fraction at higher masses. On the other hand, clusters show a significant rise whereas other environments display a moderate increase in the red fraction at lower masses. The quenching in lower mass galaxies is more influenced by their geometric environment compared to the higher mass galaxies.
(iii) We study the relative fractions of red and blue galaxies across different cosmic web environments. Filaments and sheets show a decrease in the relative red fraction (RRF) and an increase in the relative blue fraction (RBF) with decreasing redshift, while clusters exhibit the opposite trend. For lower mass galaxies, the trends in RRF and RBF are consistent with the entire sample. For higher mass galaxies, clusters initially have higher RRF and RBF than filaments. This trend reverses around $z\sim 1$. The differences in RRF and RBF across different geometric environments are primarily decided by the varying proportions of galaxies in these environments.
(iv) We investigate how the RRF and RBF of galaxies change with stellar mass and redshift in various cosmic web environments. Initially, the RRF is higher for lower mass galaxies in filaments compared to those in clusters. However, as redshift decreases, the RRF in clusters gradually surpasses that in filaments. The RRF in clusters at lower masses increases with decreasing redshift, making clusters dominant for low-mass red galaxies and filaments for intermediate masses. The RRF for clusters dominates at higher masses at all redshifts. It does not change significantly across the redshift range $3-0$, implying that massive ($\log(\frac{M_{*}}{M_{\odot}})>11$) red galaxies are mostly found in galaxy clusters. At lower masses, the RRF in clusters remains between $\sim(30-40)\%$ in this redshift range. On the other hand, RBF in filaments consistently have a higher relative proportion ($\sim(50-60)\%$) of blue galaxies, particularly at lower masses. This suggests that they host a diverse galaxy population, most likely due to less intense environmental quenching. The clusters, despite being densest, exhibit highest RBF at higher masses, indicating potential rejuvenation of star formation through interactions or mergers.
(v) The $(u-r)$ colour distribution of galaxies is unimodal at higher redshift and gradually shifts towards redder values over time, eventually showing bimodality by $z=2$, with the red peak growing most in clusters. At $z<1$, clusters have the highest median colour followed by filaments, sheets, and voids. The median colours of massive galaxies do not differ significantly across various environments in this redshift range implying that colour evolution is largely driven by mass rather than environment.
(vi) Clusters exhibit highest median mass followed by filaments, sheets and voids at all redshifts. The median stellar mass of galaxies increases with decreasing redshift across all environments. We observe a slower growth of median mass for lower mass galaxies in clusters at $z<1$. In contrast, the median mass for higher mass galaxies remain stable during $z=3-1.5$ and gradually increases at $z<1$ in all environments. The median SFR declines with decreasing redshift in all environments, with clusters dominating until $z=1.5$. At $z<1.5$, the median SFR in clusters levels out across environments and then decreases relative to others. Notably, the median SFR for massive galaxies drops to zero in all environments at $z<1.5$, indicating suppression of star formation in high-mass galaxies. The median sSFR decreases in all environments with decreasing redshifts with no significant differences across environments until $z=1.5$. At $z<1.5$, clusters exhibit the lowest median sSFR and voids and sheets exhibiting the highest values. The higher mass galaxies show larger variations in sSFR across environments.
Our study suggests that both stellar mass and geometric environments play an important role in galaxy evolution. The higher density regions host a larger fraction of blue galaxies during $z=3-2$, implying they provide a more favourable environment for star formation. However, the environmental factors start to dominate at $z<2$ reversing this trend. We find that at $z<1$, most of the massive galaxies ($\log(\frac{M_{*}}{M_{\odot}})>10.5$) are quenched irrespective of their geometric environments. This is consistent with AGN feedback model in the IllustrisTNG simulations which suggests that AGN feedback is highly effective in massive central galaxies ($\log(\frac{M_{*}}{M_{\odot}})>10.5$) [128]. Kinetic feedback from super-massive black holes (SMBHs) is the key mechanism driving the quenching of star formation in massive galaxies [129]. The fraction of quenched galaxies in TNG300 is in the range $(30-80)\%$ percent in the stellar mass range $10^{10.5}-10^{11.5}M_{\odot}$ [127]. Our results (Figure 3, Figure 4) are consistent with these findings. So the AGN feedback may have played a dominant role in quenching the high mass galaxies in all environments. However, the AGN feedback model in IllustrisTNG shows reduced effectiveness in quenching galaxies of intermediate and lower masses [128]. At lower masses ($\log(\frac{M_{*}}{M_{\odot}})<10.5$), quenching is largely controlled by environment. The fraction of red galaxies at lower masses increases with decreasing redshift at each geometric environment. At lower masses, clusters host the highest fraction of red galaxies followed by filaments, sheets and voids. Low mass galaxies often undergo environmental quenching before they become part of their current host group. The pre-processing can involve interactions with other galaxies or the effects of the surrounding medium, leading to a reduction in their gas reservoirs and star formation rates.
The relative fractions of red and blue galaxies (RRF and RBF) are influenced by both stellar mass and geometric environment. Generally, RRF increases and RBF decreases in clusters as redshift decreases. In contrast, filaments and sheets exhibit the opposite trend. At higher masses, both RRF and RBF in clusters decline with decreasing redshift, while in filaments, both show a steady increase. This anomalous behavior can be largely attributed to a growing relative abundance of intermediate and high-mass galaxies in filaments. Interestingly, clusters, despite being the densest environments, display the highest RRF and RBF for galaxies with masses $\log(\frac{M_{*}}{M_{\odot}})>11$. This suggests that massive galaxies are predominantly located in clusters, where most are quenched, although some may be rejuvenated through interactions or mergers. In contrast, filaments host a more diverse galaxy population as evidenced by the higher RBF and RRF at lower masses ($\log(\frac{M_{*}}{M_{\odot}}<11$). At lower masses, filaments exhibit relative red fractions (RRF) comparable to clusters, yet host nearly $60\%$ of blue galaxies. This suggests that the milder environmental quenching in filaments enables galaxies to undergo a wider range of evolutionary stages. A recent study [90] using the SIMBA simulation [91] suggest that shock heating in filaments can inhibit star formation. Another study [92] with the IllustrisTNG simulation finds that satellite galaxies are especially susceptible to such quenching. Our results are consistent with these findings. The low-mass quenched galaxies in filaments may represent satellite populations, while the star-forming galaxies may correspond to central galaxies with intermediate masses.
Our study reveals several key trends in galaxy evolution across different geometric environments over time. At higher redshifts, the $(u-r)$ colour distribution is unimodal but turns into bimodal by $z=2$, particularly in clusters, which have the highest median colour at $z<1$. The median colour and stellar mass of galaxies increases steadily from $z=3$ ato $z=0$. For massive galaxies, a steady increase in median colour across all environments, with little variations across environments, suggests that the process driving the reddening of massive galaxies is fairly uniform across the cosmic web. This could indicate that the quenching mechanisms, such as AGN feedback, which cause galaxies to evolve from blue to red, operate similarly across different environments. In contrast, the significant differences in median stellar mass across environments suggest that the growth of galaxy mass is more sensitive to the local cosmic web environment, with different environments supporting different rates of stellar mass buildup. This might reflect environmental factors influencing galaxy mergers or accretion more strongly than colour evolution. This is further supported by the decline in SFR and sSFR across all environments, with massive galaxies showing a sharp reduction in these rates at lower redshifts, regardless of their environment.
We also compare our results with the earlier studies on galaxy evolution with hydrodynamical simulations. [109] investigate the evolution of galaxy colour using the EAGLE simulation [130] and find that the red sequence forms around $z\sim 1$ due to the quenching of low-mass satellite galaxies and AGN feedback in more massive central galaxies. [111] demonstrate using TNG100 and TNG300 that between $70\%$ and $90\%$ of central and satellite galaxies with masses greater than or equal to $\log(\frac{M_{*}}{M_{\odot}})>10$ are predicted to be quenched in the redshift range $0\leq z\leq 0.5$. [110] demonstrate that in the EAGLE simulation, the quenching time scales for both central and satellite galaxies decrease with increasing stellar mass. They observe an increase in the fraction of passive galaxies at lower redshifts. Another study [112] using the TNG simulation finds that slow-quenching galaxies are approximately twice as common as fast-quenching galaxies in the redshift range of $20.7<z<$ 2. More recently, [113] study the roles of local density and interactions on the evolution of red and blue galaxies in the redshift range $0\leq z\leq 3$ using the EAGLE simulation. They report that red galaxies tend to occupy denser regions and show significantly stronger clustering compared to blue galaxies at $z<2$. They also find a substantial decrease in the blue fraction in paired galaxies at $z<1$ suggesting significant roles of interactions in galaxy evolution. Our results are consistent with these findings. Further, Our findings also align with the observational studies [99, 97, 131, 98], reinforcing the effectiveness of hydrodynamical simulations as a valuable tool for investigating galaxy formation and evolution.
Our study indicates that quenching at higher masses is mainly driven by stellar mass, whereas quenching in intermediate- and low-mass galaxies is influenced by a complex interplay between stellar mass and environment. We conclude that both stellar mass and cosmic web environments play a significant role in shaping colour bimodality and driving overall galaxy evolution.
ACKNOWLEDGEMENTS
BP would like to acknowledge IUCAA, Pune, for providing support through the associateship programme. AN acknowledges the financial support from the Department of Science and Technology (DST), Government of India through an INSPIRE fellowship. AN thanks Dylan Nelson for help in downloading and understanding the data from the IllustrisTNG database.
The IllustrisTNG simulations were undertaken with compute time awarded by the Gauss Centre for Supercomputing (GCS) under GCS Large-Scale Projects GCS-ILLU and GCS-DWAR on the GCS share of the supercomputer Hazel Hen at the High Performance Computing Center Stuttgart (HLRS), as well as on the machines of the Max Planck Computing and Data Facility (MPCDF) in Garching, Germany.
Data availability
The data for IllustrisTNG simulations are publicly available at https://www.tng-project.org/data/. The data produced in this work will be shared on reasonable request to the authors.

### v2

@@
 Despite being the densest environment, clusters display the highest relative blue fraction (RBF) for high-mass galaxies, likely due to interactions or mergers that can temporarily rejuvenate star formation in some of them.
-The $(u-r)$ colour distribution transitions from unimodal to bimodal by redshift $z=2$ across all environments.
+The (ur)(u-r) colour distribution transitions from unimodal to bimodal by redshift $z=2$ across all environments.
 At $z<1$, clusters exhibit the highest median colour, with stellar mass being the primary driver of colour evolution in massive galaxies.
@@
 All simulations begin at a redshift of $z=127$, using cosmological parameters from the Planck 2015 results [117]: $\Omega_{\Lambda}=0.6911$, $\Omega_{m}=0.3089$, $\Omega_{b}=0.0486$, $\sigma_{8}=0.8159$, $n_{s}=0.9667$, and $h=0.6774$.
-For our analysis, we utilize the TNG300-1 run, which offers the largest volume and highest resolution in the TNG series.
+For our analysis, we utilize the TNG300-1 run, which offers the highest resolution run among the largest available volume in the TNG series.
 The specifics of TNG300-1 are detailed in the  Table 1, and our study covers redshifts from $z=3$ to the present ($z=0$).
@@
 In this work, we select only galaxies with a non-zero SubhaloFlag, as a value of zero indicates a subhalo of non-cosmological origin.
-Additionally, we impose a stellar mass criterion, choosing galaxies in the range $129<\log_{10}(\frac{M_{\star}}{M_{\odot}})<$ 12.
+Additionally, we impose a stellar mass criterion, choosing galaxies in the range $9<\log_{10}(\frac{M_{\star}}{M_{\odot}})<12$.
 Stellar mass is defined as the total mass of star particles within twice the stellar half-mass radius $(r_{stars},1/2)$ [101].
-For $(u-r)$ colour, we use the supplementary catalog provided by  [102].
-The star formation rates are obtained from the SubhaloSFRinRad field in the group catalog.
+For (ur)(u-r) colour, we use the supplementary catalog provided by [98], which provides apparent magnitudes for different snapshots and dust-corrected absolute magnitudes only at $z=0$.
+At $z=0$, we use the dust-corrected absolute magnitudes to compute colours.
+For all other redshifts ($z>0$), we compute colours using the apparent magnitudes, as dust-corrected values are not available.
+Since colour is determined by the difference between magnitudes in two bands, and the distance modulus is the same for both bands of a given galaxy, the colour remains a relative quantity.
+Although the absence of dust correction at $z>0$ may introduce systematic offsets, our classification method is designed to be robust to such effects.
+Specifically, we apply Otsu’s thresholding method [111, 110] to the (ur)(u-r) colour distribution at each redshift and in different stellar mass bins.
+This data-driven method identifies the optimal colour threshold based on the bimodal distribution of galaxy colours, allowing for a consistent and redshift-dependent separation of red and blue populations without relying on fixed colour cuts or rest-frame corrections.
+The bimodal distribution is more prominent in (ur)(u-r) than other colours, such as (gr)(g-r).
+The u-band flux is strongly affected by young, hot stars, while r-band flux traces older stars.
+This makes (ur)(u-r) colour a better indicator of recent star formation activity, whereas (gr)(g-r) is less sensitive to changes in young stellar populations.
+We obtain the star formation rates from the SubhaloSFRinRad field in the group catalog.
 In IllustrisTNG, star formation is modeled following [120].
 For our analysis, we use the instantaneous star formation rate, which is the sum of the star formation rates of all star-forming gas cells within twice the stellar half-mass radius of the galaxy.
-Figure 1: This figure shows the distribution of galaxies in the $(u-r)$ colour-stellar mass plane across each redshift, with the dividing line (determined by Otsu’s method) distinguishing the red and blue populations.
+Figure 1: This figure shows the distribution of galaxies in the (ur)(u-r) colour-stellar mass plane across each redshift, with the dividing line (determined by Otsu’s method) distinguishing the red and blue populations.
+In each subplot, we also show $10$ uniformly spaced contours lying between the minimum and maximum density of galaxies in the respective colour-stellar mass plane.
 3 Method
 The primary objective of this study is to classify galaxies in the IllustrisTNG simulation into red and blue populations according to their stellar mass and redshift, and to investigate how cosmic web environments and stellar mass influence their evolution.
+The (ur(u-r) colour distribution is bimodal at low redshift, with peaks corresponding to the blue cloud and the red sequence.
+Traditionally a fixed (ur)(u-r) colour cut is used in many earlier studies to separate the red and blue galaxies.
+But a fixed colour threshold assumes the same division between red and blue galaxies regardless of redshift or stellar mass, which does not reflect galaxy evolution.
+The precise location of the peaks of the two populations evolves with redshift and stellar mass.
+Thus, a static threshold could misclassify galaxies, particularly at higher redshifts where the colour distribution is less bimodal.
+Otsu’s method [111, 110] is better than a fixed colour cut because it provides adaptive, optimal separation of red and blue galaxies while accounting for variations across different stellar mass bins and redshifts.
 3.1 Identifying red and blue Galaxies with Otsu’s method
@@
 This method iterates through all possible thresholds, selecting the one that provides the optimal separation between the two classes.
-This technique is particularly useful when the data distribution is bimodal, as is the case with the $(u-r)$ colour distribution of galaxies, where we observe two peaks corresponding to the blue cloud and the red sequence.
+This technique is particularly useful when the data distribution is bimodal, as is the case with the (ur)(u-r) colour distribution of galaxies, where we observe two peaks corresponding to the blue cloud and the red sequence.
 Recently, [110] proposed a method for classifying galaxies into two groups: the “blue cloud” (younger, star-forming galaxies) and the “red sequence” (older, passive galaxies) based on the Otsu’s technique.
@@
   1. 1.
-    Histogram Calculation: We first calculate the histogram of the $(u-r)$ colour for all galaxies in our sample, using a specific number of bins $M$.
+    Histogram Calculation: We first calculate the histogram of the (ur)(u-r) colour for all galaxies in our sample, using a specific number of bins $M$.
 The histogram is then normalized by the total number of galaxies $N=\sum_{i=1}^{M}n_{i}$ where $n_{i}$ corresponds to the number of galaxies in the $i^{th}$ colour bin.
-This ensures that the sum of the probabilities for all bins equals to one i.e $\sum_{i=1}^{M}p_{i}=1$ where $p_{i}=\frac{n_{i}}{N}$.
+This ensures that the total probability across all bins sums to one i.e.
+$\sum_{i=1}^{M}p_{i}=1$ where $p_{i}=\frac{n_{i}}{N}$.
   2. 2.
-    Class Probabilities: For a given threshold, the galaxies are divided into two groups: blue cloud (BC) and red sequence (RS).
-The probability of each class occurring is calculated based on the proportion of galaxies in each group relative to the total sample.
-If the threshold is at the k-th bin, the blue cloud is represented by the first $k$ bins, while the red sequence is represented by the bins from $k+1$ to the end.
-The class probabilities are given by,
+    Class Probabilities: A threshold ($k$-th bin) is chosen, dividing galaxies into blue cloud (bins $1$ to $k$) and red sequence (bins $k+1$ to $M$).
+We iterate through all possible thresholds (i.e.
+$k$ assumes all values between $1$ to M1M-1).
+The probability of a galaxy belonging to each class is computed.
+The class probabilities of the blue cloud (BC) and the red sequence (RS) are respectively given by,
     $P_{BC}=\sum_{i=1}^{k}p_{i}=w(k)$ (3.1)
@@
     $\mu_{RS}=\frac{\sum_{i=k+1}^{M}x_{i}p_{i}}{P_{RS}}=\frac{\mu_{T}-\mu_{k}}{1-w(k)}$ (3.4)
-    Here, $x_{i}$ represents the $(u-r)$ colour corresponding to the $i^{th}$ bin.
+    Here, $x_{i}$ represents the (ur)(u-r) colour corresponding to the $i^{th}$ bin.
 The mean up to the $k^{th}$ bin is given by $\mu_{k}=\sum_{i=1}^{k}x_{i}p_{i}$, while $\mu_{T}=\sum_{i=1}^{M}x_{i}p_{i}$ denotes the mean of the entire distribution.
@@
   4. 4.
-    Intra-class variance and Inter-class variance: The intra-class (within-class) variance ($\sigma_{wc}^{2}$) and the inter-class (between-class) variance ($\sigma_{bc}^{2}$) are calculated as,
+    Intra-class variance and Inter-class variance: The total variance is divided into intra-class (within-class) variance and inter-class (between-class) variance.
+The intra-class variance ($\sigma_{wc}^{2}$) and the inter-class variance ($\sigma_{bc}^{2}$) are calculated as,
     $\sigma_{wc}^{2}=P_{BC}\,\sigma_{BC}^{2}+P_{RS}\,\sigma_{RS}^{2}$ (3.7)
@@
     Threshold Selection: The optimal threshold is the one that minimizes the intra-class variance $\sigma_{wc}^{2}$, which is equivalent to maximizing the inter-class variance $\sigma_{bc}^{2}$.
-In practice, this threshold can be found by iterating through the possible values and selecting the one that satisfies these conditions.
-This yields a threshold that is not affected by binning choices, making it a reliable approach for classifying red and blue galaxies [114].
-This method offers a robust and parameter-free way to classify galaxies into the blue cloud and red sequence based on their $(u-r)$ colour and stellar mass.
+We iterate through all possible thresholds, selecting the one that minimizes the intra-class variance while maximizing the inter-class variance.
+This threshold is not affected by binning choices, making it a reliable approach for classifying red and blue galaxies [110].
 Redshift Number of galaxies
-All $9<\log_{10}\left(\frac{M_{\star}}{M_{\odot}}\right)\leq 10.5$ $1210.5<\log_{10}\left(\frac{M_{\star}}{M_{\odot}}\right)<$ 12
+All $9<\log_{10}\left(\frac{M_{\star}}{M_{\odot}}\right)\leq 10.5$ $10.5<\log_{10}\left(\frac{M_{\star}}{M_{\odot}}\right)<12$
 Total Red Blue Total Red Blue Total Red Blue
@@
 $0.7$ $242859$ $81067$ $161792$ $218375$ $64601$ $153774$ $24484$ $16466$ $8018$
-$1$ $231933$ $73719$ $158214$ $210682$ $60488$ $150194$ $21251$ $13231$ $8020$
+11 $231933$ $73719$ $158214$ $210682$ $60488$ $150194$ $21251$ $13231$ $8020$
 $1.5$ $199307$ $69796$ $129511$ $184161$ $61093$ $123068$ $15146$ $8703$ $6443$
-$2$ $157394$ $66878$ $90516$ $147377$ $61740$ $85637$ $10017$ $5138$ $4879$
-$3$ $84438$ $46230$ $38208$ $80992$ $44559$ $36433$ $3446$ $1671$ $1775$
-Table 2: This table shows the total number of galaxies with nonzero SubhaloFlag and stellar masses within $129<\log_{10}(\frac{M_{\star}}{M_{\odot}})<$ 12 in IllustrisTNG simulation for each snapshot from redshift 00 to $3$.
+22 $157394$ $66878$ $90516$ $147377$ $61740$ $85637$ $10017$ $5138$ $4879$
+33 $84438$ $46230$ $38208$ $80992$ $44559$ $36433$ $3446$ $1671$ $1775$
+Table 2: This table shows the total number of galaxies with nonzero SubhaloFlag and stellar masses within $9<\log_{10}(\frac{M_{\star}}{M_{\odot}})<12$ in IllustrisTNG simulation for each snapshot from redshift 0 to $3$.
 It also lists the number of red and blue galaxies identified with the Otsu’s method at each redshift in different mass ranges.
@@
 Therefore, a single colour threshold cannot be justified for galaxies of varying masses and luminosities.
-We divide the entire sample into several independent stellar mass bins and apply this technique to each of these bins separately.
+We divide the entire sample into several independent stellar mass bins and apply this technique to each of these bins separately, allowing the classification to adapt to how colour varies with mass.
 This provides us a dividing line between the two populations in the colour-stellar mass plane at each redshift (Figure 1).
-We also tabulate the number of red and blue galaxies in two different mass ranges between redshift $3$ to 00 in Table 2.
+Otsu’s method determines the optimal division between red and blue population at each redshift by minimizing the intra-class variance and maximizing the inter-class variance, ensuring the separation is statistically robust.
+We tabulate the number of red and blue galaxies in two different mass ranges between redshift $3$ to 0 in Table 2.
+It may be noted that at higher redshifts, particularly at $z=3$, the (ur)(u-r) colour distribution becomes less clearly bimodal (bottom right panel of Figure 11).
+One-dimensional histograms at this epoch show a largely unimodal shape, making the identification of distinct red and blue populations less straightforward.
+However, when viewed in the stellar mass-colour plane, a subtle population of massive galaxies with relatively higher (ur)(u-r) colours emerges, suggesting that a weak form of bimodality may already be developing (bottom right panel of Figure 1).
+To maintain methodological consistency across redshifts, we continue to apply Otsu’s thresholding method at $z=3$ as a data-driven way to divide the galaxy population into two relative colour classes.
+At this redshift, the resulting groups are better interpreted as “less blue” and “more blue” galaxies, rather than physically distinct red and blue populations.
+This functional classification still provides meaningful insights into environmental trends at early times.
+This approach allows for a uniform comparative analysis across cosmic time, while acknowledging the evolving nature of galaxy colour distributions.
 Figure 2: This figure displays the distribution of galaxies in various cosmic web environments at $z=0$, from a 20 Mpc thick slice of the IllustrisTNG simulation.
+The cosmic web classification is based on the eigenvalues of the deformation tensor, estimated from the density field smoothed with a Gaussian filter of width 4 Mpc.
 3.2 Identifying different morphological environments of the cosmic web
@@
 This method relies on the eigenvalues and eigenvectors of the deformation tensor to identify voids, sheets, filaments, and clusters.
-The deformation tensor $T_{ij}$ is derived from the Hessian matrix of the gravitational potential field $\Phi$, defined as:
+The deformation tensor $T_{ij}$ is derived from the Hessian matrix of the gravitational potential field $\Phi$, defined as
 $T_{ij}=\frac{\partial^{2}\Phi}{\partial x_{i}\partial x_{j}}$ (3.10)
@@
 This field is smoothed using an isotropic Gaussian filter with a width of 4 Mpc.
+The mean intergalactic separation of our samples varies between $4.79$ Mpc (at $z=0$) to $6.91$ Mpc (at $z=3$).
+Further, we have calculated the potential using a discretized density field with a grid spacing of $\sim 1.18$ Mpc.
+The smoothing scale should be larger than the grid spacing to ensure meaningful results.
+We are primarily concerned about resolving large-scale structures of the cosmic web.
+So, we have chosen a filter width of 4 Mpc for our present analysis.
+We also study the effects of varying the filter width in subsection A.2.
 Next, we compute the Fourier transform of the gravitational potential:
@@
 Each type represents a distinct geometric environment in the cosmic web, with voids being underdense and clusters representing the densest regions.
+It may be worthwhile to mention that the density field used in this study is based on the spatial distribution of galaxies, rather than the underlying dark matter field.
+While the large-scale structure is fundamentally shaped by the dark matter distribution, and galaxies are biased tracers of that field, our choice reflects the practical constraints of observational studies.
+Since dark matter is not directly observable, galaxy-based density fields provide a more realistic and consistent basis for comparison with observational data.
+Our focus on galaxy properties motivates the use of a density measure derived from the same population, facilitating a more direct link between environment and observable galaxy characteristics.
+We utilize a subset of subhalos in each redshift applying the criteria discussed in section 2.
+The total number of subhalos in each redshift are presented in the first sub-column (“Total”) of the first column (“All”) of Table 2.
+Redshift Number of Galaxies
+Total Void Sheet Filament Cluster
+0 253140 1742 40510 128945 81943
+0.1 253212 1746 40213 129829 81424
+0.2 252546 1728 39965 130091 80762
+0.3 251329 1729 39453 129785 80362
+0.4 249659 1749 38779 129826 79305
+0.5 247760 1725 38092 128924 79019
+0.7 242859 1668 36612 127139 77440
+1 231933 1580 34128 121051 75174
+1.5 199307 1187 27280 103015 67825
+2 157394 850 19497 80358 56689
+3 84438 224 7860 41229 35125
+Table 3: This table represents the number of galaxies identified in the different regions of the cosmic web at different redshifts.
+[…]
+Figure 4: This figure shows the distribution of galaxy local densities across different cosmic web environments at each redshift.
+Here, local density refers to the density estimated using the CIC interpolation scheme applied to the galaxy distribution.
+We list the number of galaxies present in each type of cosmic web environments at different redshifts in Table 3.
+We also show the evolution of the relative fraction of galaxies in each type of cosmic web environments for the entire sample and the two distinct mass bins in Figure 3.
+The results are obtained for a smoothing length of 4 Mpc.
+The left panel of Figure 3 shows that filaments generally have the highest galaxy abundance, followed by clusters, sheets, and voids across the entire redshift range.
+The proportion of galaxies in filaments and sheets increases with decreasing redshift, while the proportion in clusters decreases.
+These trends are also evident in the lower mass bin, as depicted in the middle panel of Figure 3.
+In the high mass bin, the proportion of galaxies in filaments and clusters evolves differently, as shown in the right panel of Figure 3.
+There is a crossover around $z\sim 1$, after which filaments have a higher proportion of galaxies compared to clusters.
+Prior to $z=1$, clusters dominate in terms of galaxy proportion.
+An earlier analysis [78] using the Millennium simulation [123] find that in the present Universe, the mass fraction in filaments and sheets are $\sim 50\%$ and $\sim 24\%$ respectively.
+Another study [124] with EAGLE simulation [125] find that $\sim 62-66\%$ galaxies with stellar mass larger than 5 times $10^{8}\,M_{\odot}/h$ are in filaments at $z=0$.
+Despite differences in the method, sample and simulation, it is interesting to note that our analysis indicates $\sim 52\%$ galaxies are in filaments and $\sim 18\%$ galaxies are in sheets at $z=0$.
+We also show the probability distribution function (PDF) of the local density across various cosmic web environments at different redshifts in different panels of Figure 4.
+The peaks of the distributions for different cosmic web environments occur at different densities.
+The peak for nodes appears at the highest density, followed in order by filaments, sheets, and voids.
+At each redshift, the low density regions are dominated by voids followed by sheets and filaments.
+The peak for the nodes exhibit a gradual shift towards higher densities with decreasing redshift.
 4 Results
+[…]
+It may be noted that the red fraction (RB) and blue fraction (BF) are complementary to each other, meaning that their sum always equals 1.
+However, this relationship does not extend to the relative red fraction (RRF) and relative blue fraction (RBF).
+In other words, knowing the RRF does not allow us to directly infer the RBF.
 4.1 Evolution of red and blue fractions in different cosmic web environments
 In this subsection, we examine the evolution of the fraction of red and blue galaxies in various environments within the cosmic web.
-The red fraction (RF) and blue fraction (BF) in any cosmic web environment are respectively defined as $RF=\frac{n_{R}}{n_{R}+n_{B}}$ and $BF=\frac{n_{B}}{n_{R}+n_{B}}$.
-Here, $n_{R}$ and $n_{B}$ are the number of red and blue galaxies in the same environment.
-The top left and bottom left panels of Figure 3 illustrate these fractions across different cosmic web environments.
+The top left and bottom left panels of Figure 5 illustrate the red and blue fractions (RF and BF) across different cosmic web environments.
 Although the blue fraction can be obtained by simply subtracting the red fraction from one, they are separately shown here for the sake of completeness.
-At redshift $z=3$, the blue fraction is highest in clusters, followed by filaments, sheets, and voids.
-Specifically, the blue fraction is $\sim 48\%$ in clusters and around $40\%$ in voids.
+From the bottom left panel of Figure 5, we observe that at redshift $z=3$, the blue fraction is highest in clusters, followed by filaments, sheets, and voids.
+Specifically, the blue fraction is $\sim 48\%$ in clusters and around $35\%$ in voids.
 Although the differences in blue fraction across environments at $z=3$ are not substantial, the trend suggests that overdense regions were more conducive to star formation in earlier times.
 This may arise due to the presence of larger gas reservoirs in these regions.
-As redshift decreases, the blue fraction increases in all cosmic web environments.
+The first column of Figure 5 shows that as redshift decreases, the blue fraction increases in all cosmic web environments.
 However, this trend reverses at lower redshifts, with the blue fraction declining and the red fraction rising.
@@
 These results indicate an environment-dependent trend in galaxy colour transformation.
-Physical mechanisms that suppress star formation are more dominant in clusters and become effective earlier.
-This may be due to gradual gas depletion from ongoing star formation and faster gas exhaustion from interaction-induced starbursts in high-density regions.
-Suppression of star formation is less effective in filaments and least effective in sheets and voids, as evidenced by a high blue fraction ($\sim 68\%$) in these environments.
-This is also supported by a relatively smaller reduction in the blue fraction in sheets and voids from their respective peak values ($\sim 78\%$ at $z=0.5$).
+At $z=3$, clusters exhibit the highest blue galaxy fraction, followed by filaments, sheets, and voids.
+This trend suggests that early clusters may have hosted more star-forming galaxies, potentially due to higher gas availability.
+The blue fraction increases across all environments until redshift $z=1.5$, which likely corresponds to the peak of cosmic star formation activity.
+After $z=1.5$, the blue fraction begins to decline, with red galaxies becoming dominant in clusters.
+This shift may reflect the onset of processes that suppress star formation, such as gas depletion, AGN feedback, or galaxy interactions, which are expected to be more prevalent in dense environments.
+The timing of this transition varies across environments, with filaments and sheets showing later changes compared to clusters.
+This pattern suggests that the influence of environment on galaxy evolution may depend on local density, although we caution that our analysis does not establish causality.
+The blue fraction remains highest in voids until $z<0.5$, after which an increase in red galaxy fraction is observed.
+These trends are consistent with the idea that quenching processes become effective earlier in denser environments.
+The inflection points seen in Figure 5 thus reflect a sequence in galaxy evolution that aligns with increasing environmental density, though we emphasize that these are observed trends and not direct evidence of specific physical mechanisms.
 […]
 […]
-Figure 5: Same as Figure 4 but for blue fraction.
-We now divide our galaxy sample at each redshift into two distinct mass bins: $9<\log(\frac{M_{*}}{M_{\odot}})\leq 10.5$ and $1210.5<\log(\frac{M_{*}}{M_{\odot}})<$ 12 to examine the evolution of red and blue fractions in lower and higher mass galaxies separately.
-Results for these two mass bins are presented in the middle and right panels of Figure 3.
-The middle panels display the evolution of red and blue fractions for the lower mass bin, which includes approximately $\sim 90\%$ galaxies of the entire sample (Table 2).
-The results for the lower mass bin are quite similar to those for the entire sample for this reason.
-The results for the higher mass bin, containing about $\sim 10\%$ of the galaxies, are shown in the right panels of Figure 3.
-Even for the galaxies in the higher mass bin at $z=3$, we observe a higher blue fraction and a lower red fraction in clusters compared to filaments and sheets.
+Figure 7: Same as Figure 6 but for red fraction.
+We now divide our galaxy sample at each redshift into two distinct mass bins: $9<\log(\frac{M_{*}}{M_{\odot}})\leq 10.5$ and $10.5<\log(\frac{M_{*}}{M_{\odot}})<12$ to examine the evolution of red and blue fractions in lower and higher mass galaxies separately.
+Results for these two mass bins are presented in the second and third columns of Figure 5.
+The second column of Figure 5 displays the evolution of red and blue fractions for the lower mass bin, which includes approximately $\sim 90\%$ galaxies of the entire sample (Table 2).
+For this reason, the results for the lower-mass bin closely resemble those of the full sample.
+The results for the higher mass bin, containing about $\sim 10\%$ of the galaxies, are shown in the third column of Figure 5.
+The third column of Figure 5 shows that even for the galaxies in the higher mass bin at $z=3$, we observe a higher blue fraction and a lower red fraction in clusters compared to filaments and sheets.
 However, the red fraction steadily increases across all environments as redshift decreases, indicating that star formation is gradually suppressed in massive galaxies regardless of their environment.
-Notably, the blue fraction in voids increases between $z=2$ and $z=1$, suggesting that massive galaxies in voids may follow different evolutionary trajectories compared to those in other cosmic web environments.
+Interestingly, we note that massive galaxies in voids experience an increase in star formation between $z=2$ and $z=1$, unlike similarly massive galaxies in other environments.
+The sharp increase also indicates that massive galaxies in voids may follow different evolutionary trajectories compared to those in other cosmic web environments.
+The gas from the expanding void regions can significantly enhance star formation activity in galaxies residing within them [126, 127].
+The absence of the $z=3$ point in the third column of Figure 5 is due to the lack of galaxies with masses between $10.5<\log(\frac{M_{*}}{M_{\odot}})<12$ in voids at $z=3$.
 4.2 Stellar mass dependence of the red and blue fractions in different cosmic web environments and their evolution
-We also examine the red and blue fractions as functions of stellar mass within each geometric environment.
-Figure 4 and Figure 5 show these results across different redshifts.
-The bottom right panels of these figures reveal that at $z=3$, clusters generally have a higher blue fraction and a lower red fraction across nearly all masses.
-The blue fraction in filaments is lower than in clusters but higher than in sheets and voids throughout the entire mass range.
-However, the larger error bars for voids and sheets limit the statistical significance of these differences.
-Between $z=3-1.5$, we observe that in each geometric environment, the blue fraction increases significantly in lower mass galaxies ($\log(\frac{M_{*}}{M_{\odot}}\leq 10.5$).
-The blue fraction in lower mass galaxies do not exhibit a strong dependence on the geometric environment during this period.
-A clear environment-dependent trend emerges at $z<1$.
-We find that in clusters, the red fraction of lower mass galaxies increases from approximately $\sim 40\%$ to $\sim 70\%$ between redshifts 1 and 0.
-For higher mass galaxies, the red fraction rises from about $\sim 80\%$ to $\sim 95\%$ in all environments over the same period.
-Further, $\sim 40\%$ of galaxies in filaments are red, compared to around $\sim 30\%$ in sheets and voids.
-This indicates that the quenching of lower mass galaxies is strongly influenced by their environment at $z<1$, while quenching in more massive galaxies is less affected by their embedding geometric environments during the same period.
+We analyze the red and blue galaxy fractions as functions of stellar mass across different cosmic web environments, as shown in Figures 6 and 7.
+At $z=3$, clusters exhibit a higher blue fraction and lower red fraction across most mass ranges.
+Filaments show intermediate blue fractions that are lower than clusters but higher than sheets and voids.
+However, the larger uncertainties due to a smaller number of galaxies in sheets and voids reduce the statistical significance of these differences.
+From $z=3$ to $z=1.5$, the blue fraction rises significantly for low-mass galaxies ($\log(M_{\star}/M_{\odot})\leq 10.5$) in all environments, with minimal environmental dependence.
+A clear environment-dependent trend emerges at $z<1$ for low-mass galaxies, where the red fraction increases in clusters from $\sim 40\%$ at $z=1$ to $\sim 70\%$ at $z=0$.
+At $z=0$, $\sim 40\%$ low mass galaxies in filaments are red compared to $\sim 30\%$ in sheets and voids.
+In contrast, for high-mass galaxies ($\log(M_{\star}/M_{\odot})>10.5$), the red fraction increases across all environments from $\sim 80\%$ to $\sim 95\%$ suggesting that quenching in massive galaxies is largely independent of environment.
+We observe a peak in the blue fraction at $\log(M_{\star}/M_{\odot})\sim 10.5$ between $z=2$ and $z=1$ across all cosmic web environments (Figure 6).
+This indicates that intermediate-mass galaxies were the most actively star-forming population during this epoch, regardless of their location within the cosmic web.
+The presence of this peak suggests that galaxies in this mass range had sufficient gas reservoirs and moderate feedback regulation, enabling sustained star formation before the onset of widespread quenching mechanisms at lower redshifts.
+The sharp increase in the red fraction at $\log(M_{\star}/M_{\odot})\sim 10.5$ (Figure 7) marks a critical transition in galaxy evolution, where galaxies shift from being primarily star-forming to increasingly quenched.
+Below this mass threshold, environmental factors dominate quenching, while internal mechanisms like mass quenching and AGN feedback take over above it [7, 8, 128, 129, 130, 131].
 4.3 Evolution of relative red and blue fractions in different cosmic web environments
-The red and blue fractions provide the proportion of two populations in any particular environment.
-We can use the relative red fraction (RRF) and the relative blue fraction (RBF) to study the relative abundance of red and blue galaxies in various cosmic web environments.
-The RRF and RBF are respectively defined as $RRF=\frac{(n_{R})_{i}}{\sum_{i=1}^{4}(n_{R})_{i}}$ and $RBF=\frac{(n_{B})_{i}}{\sum_{i=1}^{4}(n_{B})_{i}}$, where $(n_{R})_{i}$ and $(n_{B})_{i}$ are the numbers of red and blue galaxies in $i^{th}$-type environment.
-Here $i$ runs from $1$ to $4$ covering 4-types of cosmic web environment.
 We examine the evolution of the RRF and RBF across different cosmic web environments, as shown in the top and bottom left panels of Figure 8.
-The RRF and RBF represent the proportions of red and blue galaxies, respectively, within a specific environment compared to other cosmic environments.
 The top and bottom left panels of Figure 8 reveal that in filaments and sheets, the RRF decreases and the RBF increases with decreasing redshift.
@@
 This counter-intuitive result can be explained by examining the relative proportions of galaxies in different cosmic web environments.
-Figure 7 illustrates the evolution of the relative fractions of galaxies in filaments, sheets, clusters, and voids for the entire sample and the two distinct mass bins.
-The left panel of Figure 7 shows that filaments generally have the highest galaxy abundance, followed by clusters, sheets, and voids across the entire redshift range.
-The proportion of galaxies in filaments and sheets increases with decreasing redshift, while the proportion in clusters decreases.
-These trends are also evident in the lower mass bin, as depicted in the middle panel of Figure 7.
-In the high mass bin, the proportion of galaxies in filaments and clusters evolves differently, as shown in the right panel of Figure 7.
-There is a crossover around $z\sim 1$, after which filaments have a higher proportion of galaxies compared to clusters.
-Prior to $z=1$, clusters dominate in terms of galaxy proportion.
 […]
-[…]
-The anomalous results in the right panels of Figure 6 are explained by the trends observed in the right panel of Figure 7 and Figure 3.
+The anomalous results in the right panels of Figure 8 can be explained by the trends observed in the right panel of Figure 3 and Figure 5.
 For high mass galaxies, RF increases and BF decreases in all types of cosmic web environments, with comparable values across different environments (Figure 5).
@@
 The trends observed in the right panel of Figure 3 are reflected in the top and bottom right panels of Figure 8, with crossovers in RRF and RBF occurring at different redshifts due to the varying red and blue fractions in filaments and clusters.
-It may be noted that the red fraction (RB) and blue fraction (BF) are complementary to each other, meaning that their sum always equals 1.
-However, this relationship does not extend to the relative red fraction (RRF) and relative blue fraction (RBF).
-In other words, knowing the RRF does not allow us to directly infer the RBF.
 4.4 Stellar mass dependence of the relative red and blue fractions in different cosmic web environments and their evolution
-We now examine the RRF and RBF across different cosmic web environments as functions of stellar mass and their evolution from redshift $3$ to 00.
-The various panels in Figure 8 display the RRF as a function of stellar mass in different cosmic web environments at different redshifts.
-The results indicate that in both filaments and sheets, the RRF is lower at higher masses and higher at lower masses.
-The RRF in filaments and sheets at lower masses decreases with decreasing redshift.
-Conversely, clusters exhibit an opposite trend, where the RRF is higher at higher masses and lower at lower masses.
-In clusters, the RRF at higher masses remains relatively stable with redshift, while at lower masses, it gradually increases as redshift decreases.
-At $z\sim 1$, the RRF at lower masses becomes comparable between filaments and clusters.
-However, below this redshift, clusters tend to dominate at lower masses, while filaments become more prominent at intermediate masses.
-At low redshifts, the RRF in clusters exceeds that of all other environments at both low and high masses.
-This suggests that low-mass red galaxies in clusters may represent quenched satellites, while high-mass red galaxies may be massive centrals.
-Interestingly, the RRF in clusters and filaments shows two distinct crossovers at intermediate masses.
-Between these crossovers ($1110.5<\log(\frac{M_{*}}{M_{\odot}})<$ 11), the RRF in filaments surpasses that of all other environments.
-It may be worthwhile to mention here that the right panels of Figure 6 indicate that, at lower redshifts, both RRF and RBF dominate in filaments for masses $\log(\frac{M_{*}}{M_{\odot}})>10.5$.
-However, a closer examination reveals that this dominance is confined to the mass range $1110.5<\log(\frac{M_{*}}{M_{\odot}})<$ 11.
-There are only a small number of red and blue galaxies with masses $\log(\frac{M_{*}}{M_{\odot}})>11$, and these primarily reside in clusters.
+We examine the evolution of the RRF and RBF as functions of stellar mass and cosmic web environment from $z=3$ to $z=0$ (Figure 9, Figure 10).
+In filaments and sheets, the RRF is relatively higher at low stellar masses compared to higher masses and decreases with decreasing redshift, indicating that low-mass galaxies in these environments tend to remain blue for longer periods.
+Conversely, in clusters, the RRF is initially higher at high masses but increases at low masses over time, suggesting that environmental quenching of low-mass galaxies becomes more prominent at late times.
+At $z\sim 1$, the RRF at low masses becomes comparable in filaments and clusters.
+By $z=0$, however, clusters dominate the red population at both low and high stellar masses, while filaments show a peak in RRF at intermediate masses ($10.5<\log(M_{\star}/M_{\odot})<11$).
+Two notable crossovers in the RRF curves at $z=0$ illustrate this shift: at low masses, clusters overtake filaments in RRF, indicating intensified quenching in dense environments; and at intermediate masses, filaments surpass clusters, emerging as the primary hosts of quiescent galaxies in that mass range.
+RBF trends provide a complementary view.
+Clusters consistently show lower RBF values at low stellar masses across redshifts, which is consistent with clusters hosting lower fractions of blue, low-mass galaxies.
+In contrast, filaments show the highest RBF at intermediate and low masses for $z<1$, highlighting their transitional role hosting both star-forming and quenched galaxies.
+This contrasts with clusters, which are associated with efficient quenching and consistently exhibit low RBF values compared to filaments across redshifts.
+Voids, on the other hand, contain a very small number of galaxies overall, resulting in low RRF and RBF values at all stellar masses and redshifts.
+While their low-density nature suggests a star-forming preference, the small sample size limits statistical interpretation.
+These trends support the idea that mass-driven and environment-driven quenching processes operate in tandem, but on different timescales and with varying efficiencies across environments.
+Filaments appear to quench galaxies earlier, particularly at low stellar masses, while clusters become increasingly effective at quenching low-mass galaxies at later times.
+For intermediate-mass galaxies, filaments play a dominant role significantly influencing their evolutionary pathways.
+Our results are consistent with previous findings [132, 130], which show that low-mass galaxies in dense environments quench rapidly, whereas intermediate-mass galaxies are more sensitive to the surrounding large-scale structure.
+Interestingly, clusters also exhibit the highest RBF at the high-mass end ($\log(M_{\star}/M_{\odot})>11$) even at low redshift, possibly reflecting episodes of renewed star formation triggered by mergers or dynamical interactions in these dense environments.
 […]
 Figure 10: Same as Figure 9 but for RBF.
-We present the RBF as a function of stellar mass across different cosmic web environments and redshifts in the various panels of Figure 9.
-Our analysis reveals that in both filaments and sheets, the RBF is lower at higher masses and higher at lower masses.
-Specifically, the RBF in filaments and sheets at lower masses increases as redshift decreases.
-Between redshift $z=1-0$, filaments exhibit a significantly higher RBF ($\sim 0.55$) compared to other environments for galaxies having $\log(\frac{M_{*}}{M_{\odot}})<11$.
-This indicates that filaments host a diverse population of galaxies, encompassing both passive and star-forming galaxies.
-The less aggressive environmental quenching in filaments allows for a wider range of evolutionary stages.
-Interestingly, despite being the densest environment, clusters show the highest RBF for high-mass galaxies ($\log(\frac{M_{*}}{M_{\odot}}>11$) even at lower redshifts.
-This suggests that clusters are highly dynamic environments where interactions or mergers might temporarily reignite star formation activity in some high mass galaxies.
-Figure 10: This figure shows the PDF of $(u-r)$ colour in different cosmic web environments at different redshifts.
+Figure 11: This figure shows the PDF of (ur)(u-r) colour in different cosmic web environments at different redshifts.
 […]
+The trends observed in all the galaxy fractions across cosmic web environments may, in part, be driven by variations in local density, since denser regions naturally host different types of galaxies compared to underdense areas.
+To disentangle the effects of local density from those of the geometric environment, it is essential to examine the fractions F, RF, BF, RRF, and RBF within narrow density bins.
+This approach would allow us to assess whether the observed environmental trends are intrinsic to the cosmic web environment or simply a reflection of underlying density variations, providing a clearer picture of the role played by cosmic web in galaxy evolution.
+We address this issue in detail in subsection A.1 by analyzing the galaxy fractions within narrow density bins to isolate the influence of cosmic web geometry from that of density.
 4.5 Evolution of colour across different cosmic web environments
-We present the distributions of $(u-r)$ colour across various cosmic web environments at different redshifts, ranging from $z=3$ to $z=0$, as shown in different panels of Figure 10.
-At $z=3$, the $(u-r)$ colour distribution in each environment is unimodal, with a peak around $(u-r)\sim 0.7$.
-As redshift decreases, this peak shifts towards higher $(u-r)$ values, indicating the aging of stellar populations.
+We present the distributions of (ur)(u-r) colour across various cosmic web environments at different redshifts, ranging from $z=3$ to $z=0$, as shown in different panels of Figure 11.
+At $z=3$, the (ur)(u-r) colour distribution in each environment is unimodal, with a peak around $(u-r)\sim 0.7$.
+As redshift decreases, this peak shifts towards higher (ur)(u-r) values, indicating the aging of stellar populations.
 By $z=2$, the colour distribution begins to show bimodality.
@@
 The emergence of distinct bimodality across all environments suggests that the colour transformation can occur in any cosmic web environment.
-In the top left panel of Figure 11, we show the median $(u-r)$ colour of galaxies in various cosmic web environments as a function of redshift.
+In the top left panel of Figure 12, we show the median (ur)(u-r) colour of galaxies in various cosmic web environments as a function of redshift.
 Since $z\sim 2$, the median colour in clusters begins to diverge from that in other environments and increases rapidly after $z>1$.
@@
 The lower mass bin shows trends similar to the combined sample due to the large number of galaxies in this bin.
-The bottom panel reveals that the median colour for more massive galaxies ($1210.5<\log(\frac{M_{*}}{M_{\odot}})<$ 12) increases steadily at $z<1$, with little variation across different environments.
+The bottom panel reveals that the median colour for more massive galaxies ($10.5<\log(\frac{M_{*}}{M_{\odot}})<12$) increases steadily at $z<1$, with little variation across different environments.
 This suggests that the colour evolution of the more massive galaxies is primarily driven by mass and is less dependent on their environment.
@@
 The top right and bottom panels of Figure 14 show the results for lower and higher mass bins, respectively.
-In the top right panel, we observe that the median mass of galaxies in filaments and sheets increases steadily between $z=1$ and $z=0$.
+In the top right panel, we observe that the median mass of galaxies in filaments and sheets increases steadily between $z=3$ to $z=1$.
 However, at $z<1$, the growth rate of the median mass for lower mass galaxies in clusters slows down significantly compared to those in filaments and sheets.
@@
 The median sSFR across different environments is shown in Figure 18.
-The top left panel indicates that, unlike the median SFR, there are no statistically significant differences in the median sSFR across environments in the redshift range $31.5<z<$ 3.
+The top left panel indicates that, unlike the median SFR, there are no statistically significant differences in the median sSFR across environments in the redshift range $1.5<z<3$.
 As redshift decreases, sSFR decreases in all environments.
@@
 We study the evolution of red and blue galaxies in different cosmic web environments using IllustrisTNG simulation. Our key findings are summarized as follows.
-(i) We track the evolution of the red and blue fractions across different cosmic web environments from redshift $z=3$ to $z=0$.
-Initially, blue galaxies are more prevalent in clusters compared to other geometric environments.
-This trend reverses at lower redshifts.
-The red fraction starts to rise after specific redshifts for each environment, with clusters experiencing earlier suppression of star formation compared to filaments, sheets, and voids.
-The massive galaxies show a consistent increase in red fractions over time in all environments.
-(ii) We analyze the red and blue fractions as a function of stellar mass across different environments at various redshifts.
-At $z=3$, clusters have a higher blue fraction and lower red fraction compared to other environments at all masses, with filaments showing an intermediate blue fraction.
-Over time, the red fraction increases in higher mass galaxies and decreases in lower mass galaxies at all environments until $z=1.5$.
-A pronounced environment-dependent trend emerges at $z<1$ during which all environments show a significant increase in the red fraction at higher masses.
-On the other hand, clusters show a significant rise whereas other environments display a moderate increase in the red fraction at lower masses.
-The quenching in lower mass galaxies is more influenced by their geometric environment compared to the higher mass galaxies.
-(iii) We study the relative fractions of red and blue galaxies across different cosmic web environments.
-Filaments and sheets show a decrease in the relative red fraction (RRF) and an increase in the relative blue fraction (RBF) with decreasing redshift, while clusters exhibit the opposite trend.
-For lower mass galaxies, the trends in RRF and RBF are consistent with the entire sample.
-For higher mass galaxies, clusters initially have higher RRF and RBF than filaments.
-This trend reverses around $z\sim 1$.
-The differences in RRF and RBF across different geometric environments are primarily decided by the varying proportions of galaxies in these environments.
-(iv) We investigate how the RRF and RBF of galaxies change with stellar mass and redshift in various cosmic web environments.
-Initially, the RRF is higher for lower mass galaxies in filaments compared to those in clusters.
-However, as redshift decreases, the RRF in clusters gradually surpasses that in filaments.
-The RRF in clusters at lower masses increases with decreasing redshift, making clusters dominant for low-mass red galaxies and filaments for intermediate masses.
-The RRF for clusters dominates at higher masses at all redshifts.
-It does not change significantly across the redshift range $3-0$, implying that massive ($\log(\frac{M_{*}}{M_{\odot}})>11$) red galaxies are mostly found in galaxy clusters.
-At lower masses, the RRF in clusters remains between $\sim(30-40)\%$ in this redshift range.
-On the other hand, RBF in filaments consistently have a higher relative proportion ($\sim(50-60)\%$) of blue galaxies, particularly at lower masses.
-This suggests that they host a diverse galaxy population, most likely due to less intense environmental quenching.
-The clusters, despite being densest, exhibit highest RBF at higher masses, indicating potential rejuvenation of star formation through interactions or mergers.
-(v) The $(u-r)$ colour distribution of galaxies is unimodal at higher redshift and gradually shifts towards redder values over time, eventually showing bimodality by $z=2$, with the red peak growing most in clusters.
-At $z<1$, clusters have the highest median colour followed by filaments, sheets, and voids.
-The median colours of massive galaxies do not differ significantly across various environments in this redshift range implying that colour evolution is largely driven by mass rather than environment.
-(vi) Clusters exhibit highest median mass followed by filaments, sheets and voids at all redshifts.
-The median stellar mass of galaxies increases with decreasing redshift across all environments.
-We observe a slower growth of median mass for lower mass galaxies in clusters at $z<1$.
-In contrast, the median mass for higher mass galaxies remain stable during $z=3-1.5$ and gradually increases at $z<1$ in all environments.
-The median SFR declines with decreasing redshift in all environments, with clusters dominating until $z=1.5$.
-At $z<1.5$, the median SFR in clusters levels out across environments and then decreases relative to others.
-Notably, the median SFR for massive galaxies drops to zero in all environments at $z<1.5$, indicating suppression of star formation in high-mass galaxies.
-The median sSFR decreases in all environments with decreasing redshifts with no significant differences across environments until $z=1.5$.
-At $z<1.5$, clusters exhibit the lowest median sSFR and voids and sheets exhibiting the highest values.
-The higher mass galaxies show larger variations in sSFR across environments.
-Our study suggests that both stellar mass and geometric environments play an important role in galaxy evolution.
-The higher density regions host a larger fraction of blue galaxies during $z=3-2$, implying they provide a more favourable environment for star formation.
-However, the environmental factors start to dominate at $z<2$ reversing this trend.
-We find that at $z<1$, most of the massive galaxies ($\log(\frac{M_{*}}{M_{\odot}})>10.5$) are quenched irrespective of their geometric environments.
-This is consistent with AGN feedback model in the IllustrisTNG simulations which suggests that AGN feedback is highly effective in massive central galaxies ($\log(\frac{M_{*}}{M_{\odot}})>10.5$) [128].
-Kinetic feedback from super-massive black holes (SMBHs) is the key mechanism driving the quenching of star formation in massive galaxies [129].
-The fraction of quenched galaxies in TNG300 is in the range $(30-80)\%$ percent in the stellar mass range $10^{10.5}-10^{11.5}M_{\odot}$ [127].
-Our results (Figure 3, Figure 4) are consistent with these findings.
-So the AGN feedback may have played a dominant role in quenching the high mass galaxies in all environments.
-However, the AGN feedback model in IllustrisTNG shows reduced effectiveness in quenching galaxies of intermediate and lower masses [128].
-At lower masses ($\log(\frac{M_{*}}{M_{\odot}})<10.5$), quenching is largely controlled by environment.
-The fraction of red galaxies at lower masses increases with decreasing redshift at each geometric environment.
-At lower masses, clusters host the highest fraction of red galaxies followed by filaments, sheets and voids.
-Low mass galaxies often undergo environmental quenching before they become part of their current host group.
-The pre-processing can involve interactions with other galaxies or the effects of the surrounding medium, leading to a reduction in their gas reservoirs and star formation rates.
-The relative fractions of red and blue galaxies (RRF and RBF) are influenced by both stellar mass and geometric environment.
-Generally, RRF increases and RBF decreases in clusters as redshift decreases.
-In contrast, filaments and sheets exhibit the opposite trend.
-At higher masses, both RRF and RBF in clusters decline with decreasing redshift, while in filaments, both show a steady increase.
-This anomalous behavior can be largely attributed to a growing relative abundance of intermediate and high-mass galaxies in filaments.
-Interestingly, clusters, despite being the densest environments, display the highest RRF and RBF for galaxies with masses $\log(\frac{M_{*}}{M_{\odot}})>11$.
-This suggests that massive galaxies are predominantly located in clusters, where most are quenched, although some may be rejuvenated through interactions or mergers.
-In contrast, filaments host a more diverse galaxy population as evidenced by the higher RBF and RRF at lower masses ($\log(\frac{M_{*}}{M_{\odot}}<11$).
-At lower masses, filaments exhibit relative red fractions (RRF) comparable to clusters, yet host nearly $60\%$ of blue galaxies.
-This suggests that the milder environmental quenching in filaments enables galaxies to undergo a wider range of evolutionary stages.
-A recent study [90] using the SIMBA simulation [91] suggest that shock heating in filaments can inhibit star formation.
-Another study [92] with the IllustrisTNG simulation finds that satellite galaxies are especially susceptible to such quenching.
-Our results are consistent with these findings.
-The low-mass quenched galaxies in filaments may represent satellite populations, while the star-forming galaxies may correspond to central galaxies with intermediate masses.
-Our study reveals several key trends in galaxy evolution across different geometric environments over time.
-At higher redshifts, the $(u-r)$ colour distribution is unimodal but turns into bimodal by $z=2$, particularly in clusters, which have the highest median colour at $z<1$.
-The median colour and stellar mass of galaxies increases steadily from $z=3$ ato $z=0$.
-For massive galaxies, a steady increase in median colour across all environments, with little variations across environments, suggests that the process driving the reddening of massive galaxies is fairly uniform across the cosmic web.
-This could indicate that the quenching mechanisms, such as AGN feedback, which cause galaxies to evolve from blue to red, operate similarly across different environments.
-In contrast, the significant differences in median stellar mass across environments suggest that the growth of galaxy mass is more sensitive to the local cosmic web environment, with different environments supporting different rates of stellar mass buildup.
-This might reflect environmental factors influencing galaxy mergers or accretion more strongly than colour evolution.
-This is further supported by the decline in SFR and sSFR across all environments, with massive galaxies showing a sharp reduction in these rates at lower redshifts, regardless of their environment.
-We also compare our results with the earlier studies on galaxy evolution with hydrodynamical simulations.
-[109] investigate the evolution of galaxy colour using the EAGLE simulation [130] and find that the red sequence forms around $z\sim 1$ due to the quenching of low-mass satellite galaxies and AGN feedback in more massive central galaxies.
-[111] demonstrate using TNG100 and TNG300 that between $70\%$ and $90\%$ of central and satellite galaxies with masses greater than or equal to $\log(\frac{M_{*}}{M_{\odot}})>10$ are predicted to be quenched in the redshift range $0\leq z\leq 0.5$.
-[110] demonstrate that in the EAGLE simulation, the quenching time scales for both central and satellite galaxies decrease with increasing stellar mass.
-They observe an increase in the fraction of passive galaxies at lower redshifts.
-Another study [112] using the TNG simulation finds that slow-quenching galaxies are approximately twice as common as fast-quenching galaxies in the redshift range of $20.7<z<$ 2.
-More recently, [113] study the roles of local density and interactions on the evolution of red and blue galaxies in the redshift range $0\leq z\leq 3$ using the EAGLE simulation.
-They report that red galaxies tend to occupy denser regions and show significantly stronger clustering compared to blue galaxies at $z<2$.
-They also find a substantial decrease in the blue fraction in paired galaxies at $z<1$ suggesting significant roles of interactions in galaxy evolution.
-Our results are consistent with these findings.
-Further, Our findings also align with the observational studies [99, 97, 131, 98], reinforcing the effectiveness of hydrodynamical simulations as a valuable tool for investigating galaxy formation and evolution.
-Our study indicates that quenching at higher masses is mainly driven by stellar mass, whereas quenching in intermediate- and low-mass galaxies is influenced by a complex interplay between stellar mass and environment.
-We conclude that both stellar mass and cosmic web environments play a significant role in shaping colour bimodality and driving overall galaxy evolution.
+(i) The evolution of red and blue fractions show strong environmental dependence.
+At $z=3$, clusters host the highest fraction of blue galaxies, but this trend reverses at lower redshifts as environmental quenching becomes dominant.
+The inflection points at $z=1.5$ for clusters, $z=0.7$ for filaments and sheets, and $z=0.5$ for voids mark key transitions where red galaxies start to dominate in these environments.
+(ii) The environmental quenching plays a major role at low masses.
+For low-mass galaxies ($\log(\frac{M_{*}}{M_{\odot}})<10.5$), clusters host the highest red fractions at $z<1$, showing that environmental effects drive quenching in these systems.
+Galaxies in filaments exhibit intermediate quenching signatures, spanning a broader range of evolutionary stages.
+At $z<1$, low-mass galaxies in filaments show a higher blue fraction compared to those in clusters.
+(iii) The filaments serve as a bridge between the galaxy evolution in cluster and void type environments.
+At $z<1$, filaments host a mix of quiescent and star-forming galaxies, with nearly $60\%$ of low-mass galaxies remaining blue.
+The presence of mixed galaxy populations in filaments may reflect both the less extreme environmental conditions, potentially allowing more gradual quenching, and the fact that filaments host a large fraction of the galaxy population, thus capturing a broad diversity of evolutionary stages.
+(iv) Our findings support the interpretation that stellar mass plays a significant role in the quenching of high-mass galaxies.
+For massive galaxies ($\log(\frac{M_{*}}{M_{\odot}})>10.5$), quenching occurs independent of environment, suggesting that internal processes like mass quenching and AGN feedback dominate over external effects.
+(v) The sharp increase in the red fraction at $\log(\frac{M_{*}}{M_{\odot}})\sim 10.5$ marks a transition in galaxy evolution.
+It corresponds to a buildup of intermediate-mass galaxies and represents a shift from star-forming to quiescent populations, becoming more prominent at $z<1$.
+(vi) The colour distribution transitions from unimodal ($z>2$) to bimodal ($z<2$), with clusters developing the strongest red peak.
+The strengthening of colour bimodality over time indicates the build-up of quiescent galaxies at lower redshifts.
+(vii) The median colours of massive galaxies do not differ significantly across various environments implying that colour evolution in high mass galaxies is largely driven by mass rather than environment.
+The median mass of lower mass galaxies in clusters exhibit a slower growth at $z<1$.
+This indicates that environmental factors influence galaxy mergers or accretion more strongly than colour evolution.
+Our findings align well with previous observational and simulation-based studies on galaxy evolution within the cosmic web, each probing specific redshift regimes and scales.
+For example, studies such as [133] (SDSS, $0.02\leq z\leq 0.155$), [134] (VIPERS, $z\sim 0.7$), [135] (COSMOS, $0.5<z<0.9$), [136] (GAMA, $0.03\leq z\leq 0.25$), [137] (WISExSCOS, $0.1<z<0.3$), [138] (SAMI, $0.004<z<0.13$), and [139] (WISE x SCOSMOS, $0.1<z<0.3$) report that galaxies near filament spines tend to be redder, more massive, and less star-forming at fixed local density on scales of a few to several Mpc.
+However, as pointed out by [140] (SDSS, $z\leq 0.077$), some of these trends may be fully explained by local density effects alone, without invoking additional influence from the filament geometry.
+One widely accepted interpretation is that galaxies migrating toward filament spines experience elevated accretion rates, more frequent interactions, and increased merger activity leading to accelerated mass growth and quenching.
+This interpretation is supported by simulation-based work across various redshifts and scales, including IllustrisTNG and zoom-in hydrodynamical studies [87, 141, 142, 143, 144, 145, 146].
+Additional processes, such as pre-processing in small galaxy groups embedded within filaments [147, 148] (z=0.150.7z=0.15-0.7), and potential shock heating in filaments suppressing star formation especially in satellite galaxies are also implicated by recent analyses using SIMBA and IllustrisTNG simulations [149, 150] (z=02z=0-2).
+Our observation of an elevated relative red fraction in filaments at intermediate stellar masses may be connected to these combined mechanisms.
+We also reinforce the picture of filaments as transitional environments.
+Observational studies such as [151] and [152], spanning $0.5\lesssim z\lesssim 2$, show that low-mass galaxies in filaments maintain higher star formation rates than their cluster counterparts at fixed stellar mass and local density.
+High-resolution zoom-in simulations by [153] find that anisotropic gas accretion onto filament halos at high redshifts ($z=2.5,4$) enhances their baryon and stellar fractions relative to field analogs.
+Dark matter-only and hydrodynamical simulations [154, 124] (z=03z=0-3) similarly indicate that low-mass halos preferentially accrete perpendicularly to their host filament, while higher-mass halos accrete along the filament axis.
+Observations of HI content in SIMBA simulations [155] (z=06z=0-6) and related anisotropic inflow signatures [156] (z=02z=0-2) further reinforce this scenario.
+Finally, the transition in colour distribution from unimodal at $z>2$ to bimodal at $z<1$ supports findings from the EAGLE simulation [105, 106], which identified buildup of the red sequence around $z\sim 1$ due to AGN feedback and satellite quenching.
+The increasing red fractions at $\log(M_{*}/M_{\odot})>10.5$ are also consistent with the mass-dependent quenching trends reported in [130, 157, 131].
+Overall, our results align with the broader consensus that environment-driven quenching dominates for low-mass galaxies, while mass- and AGN-driven quenching become increasingly important at higher masses [128, 129].
+Throughout our analysis, the cosmic web classification is based on a Gaussian-smoothed galaxy density field with a smoothing scale of 4 Mpc.
+This choice is designed to probe large-scale structures such as filaments and sheets, which typically span several megaparsecs in width.
+We emphasize that our results pertain to this intermediate scale and may not be directly comparable to studies that focus on much larger-scale structures (e.g., superclusters) or smaller-scale environments (e.g., galaxy groups or cluster cores).
+Care should be taken when interpreting similarities or differences across studies that operate at significantly different spatial resolutions or employ alternative definitions of cosmic web environments.
+The cosmic web serves as the primary source of fuel for galaxy growth, with its filamentary network playing a crucial role in galaxy formation and evolution.
+Our findings emphasize the intricate interplay between galaxy mass and environment, both of which collectively shape the evolutionary pathways of galaxies within the cosmic web.
 ACKNOWLEDGEMENTS
+Authors thank an anonymous reviewer for insightful comments and suggestions that helped us to improve the draft.
 BP would like to acknowledge IUCAA, Pune, for providing support through the associateship programme.
@@
 The data for IllustrisTNG simulations are publicly available at https://www.tng-project.org/data/. The data produced in this work will be shared on reasonable request to the authors.
+Appendix A Appendix
+We examine how local density and filter width affect our results, with detailed findings for different density bins and filter widths presented in the following two subsections.
+A.1 The impact of varying local density on galaxy fractions
+Figure 19: This shows the redshift evolution of total fraction (F), red fraction (RF), blue fraction (BF), relative red fraction (RRF), and relative blue fraction (RBF) across different cosmic web environments.
+Each row corresponds to one specific fraction, while each column shows results for a narrow bin in local density, $\rho$, with density ranges labeled at the top.
+We would like to disentangle the relative influence of local density and cosmic web environment on galaxy colour evolution.
+We examine the redshift evolution of total fraction (F), red fraction (RF), blue fraction (BF), relative red fraction (RRF), and relative blue fraction (RBF) in different cosmic web environments across 5 narrow local density bins in Figure 19.
+This allows a controlled comparison of these fractions at fixed densities.
+Each row presents the redshift evolution of one fraction, while each column corresponds to a specific density range.
+Differences observed across environments within the same density bin reveal the distinct impact of the geometric environments of the cosmic web, independent of local density.
+We note that within each local density bin, clear and consistent differences persist between the cosmic web environments which indicates that cosmic web geometry influences galaxy evolution independently of local density.
+In the top row of Figure 19, which shows the total fraction (F) of galaxies in each cosmic web environment, a consistent pattern emerges across all density bins.
+In each local density bin, filaments consistently host the largest fraction of galaxies, followed by sheets, with voids and clusters contributing smaller fractions.
+As we move from low to high density bins (left to right), the fraction of galaxies in clusters gradually increases, while the fraction in sheet decreases, as expected.
+However, filaments dominate across all bins, regardless of local density.
+This suggests that filaments are the most common environment for galaxies across the full range of densities examined.
+The fraction of galaxies in clusters dominates that in filaments only at higher densities and higher redshifts.
+The second row of Figure 19 shows that RF increases over time across all environments, with the rate of increase depending on both local density and cosmic web geometry.
+In all density bins, clusters consistently exhibit the highest red fractions, followed by filaments, sheets, and voids, especially at lower redshifts.
+As local density increases, overall RF values rise, but the ordering of environments remains stable, indicating that geometric environment influences quenching beyond what local density alone accounts for.
+An exactly opposite trend is observed in the RF at higher redshifts where RF is highest in voids followed by sheets, filaments and clusters.
+These results imply that both local density and cosmic web environment shape the build-up of red galaxies, with geometry modulating quenching efficiency even at fixed density.
+The third row of the Figure 19 shows the BF, which is complementary to the RF and reflects the relative abundance of star-forming galaxies in each environment and density bin.
+These results are directly related to the results presented in the second row of Figure 19 and support the interpretation that star formation is more sustained in underdense environments at lower redshift, while it is more efficient in denser and clustered regions at higher redshifts.
+Figure 20: This shows the impact of Gaussian filter width on the redshift evolution of total fraction (F), red fraction (RF), blue fraction (BF), relative red fraction (RRF), and relative blue fraction (RBF) across different cosmic web environments.
+Each row represents one galaxy fraction, while each column corresponds to a different filter width.
+The fourth (RRF) and fifth (RBF) rows of Figure 19 reveal trends that closely mirror those in the total fraction (F) shown in the first row, and this similarity provides deeper insight into the interplay between galaxy population demographics and environmental influence within fixed local density bins.
+In all density bins, filaments consistently dominate the total galaxy population, as shown in the first row.
+Correspondingly, in both the RRF and RBF, filaments also contribute the largest relative share of both red and blue galaxies, particularly at intermediate densities and redshifts.
+This co-dominance arises not necessarily from an intrinsic tendency for filaments to favour one galaxy type, but from the sheer number of galaxies they host, positioning them as the primary channel for both ongoing star formation and quenching within the cosmic web.
+The alignment between total fraction (F), RRF and RBF highlights an important point.
+Geometric environment shapes not just galaxy properties, but also where galaxies reside, and these distributions, in turn, amplify the environment’s role in galaxy evolution.
+Filaments are central to this narrative.
+They act as both the main reservoirs of galaxies and the transitional sites where galaxies experience a mix of quenching and star formation.
+Clusters and voids represent the extremes, favouring quenching and sustained star formation respectively.
+In essence, the combined trends in RRF, RBF, and F underscore that galaxy colour evolution is not merely about local density or intrinsic processes, but is profoundly modulated by where galaxies live in the cosmic web.
+Filaments, due to their prevalence and mixed influence, play a pivotal and nuanced role in shaping galaxy populations.
+The shifting cross-overs in the first, fourth, and fifth rows show that the influence of cosmic web environments on galaxy properties evolves with local density.
+At low densities, filaments dominate in both total fraction and relative red and blue fractions, but as density increases, clusters overtake filaments in hosting red galaxies, reflecting enhanced environmental quenching.
+These trends highlight a dynamic interplay between local density and geometric environment, where density modulates but does not replace the impact of cosmic web environment.
+Figure 19 clearly demonstrates that galaxy colour evolution is not solely governed by local density, but is also significantly influenced by the cosmic web environment.
+Across all density bins, distinct differences persist in the red and blue fractions among voids, sheets, filaments, and clusters.
+These differences cannot be explained by density alone.
+Our results suggest that galaxy colour evolution is driven by both local density and cosmic web environment.
+While local density controls the overall quenching trend, the geometric environment modulates the efficiency and timing of quenching, even at fixed local density.
+A.2 The impact of varying filter width on galaxy fractions
+We study the influence of the finite width of the Gaussian filter on our results in this subsection.
+The Figure 20 shows that the filter width used in cosmic web classification significantly influences the observed galaxy fractions and their evolutionary trends.
+While the overall qualitative behaviors persist across different filter widths, where clusters dominate red fractions, voids dominate blue fractions, and filaments host the majority of galaxies.
+However the relative contributions shift with filter width.
+Smaller filter widths better resolve transitional environments like filaments, crucial for understanding gradual quenching processes.
+At smaller widths (2 Mpc), the structures in the cosmic web are more finely resolved, allowing filaments and sheets to play a prominent role in hosting both red and blue galaxies.
+In contrast, a larger filter width smooth over finer structures blurring the boundaries between high-density peaks and surrounding structures.
+As the filter width increases, clusters contribute less to the RRF and RBF.
+This indicates that clusters become less distinct in the smoothed density field, likely being diluted into surrounding filaments and sheets due to loss of high-resolution contrast.
+Filaments and sheets, in contrast, gain a larger share of both red and blue galaxies at larger filter widths These results highlight that filter width is not a neutral choice.
+It modulates the apparent dominance of different environments and thus shapes our interpretation of how galaxies evolve within the cosmic web.
+Therefore, the choice of filter width should be carefully tailored to the characteristics and resolution of the dataset being analyzed.
+Based on these findings, we adopt a Gaussian smoothing scale of 4 Mpc for our main analysis.
+This choice reflects a careful balance between spatial resolution and structural coherence.
+A 4 Mpc filter is slightly below the mean intergalactic separation in our sample ($\sim 5$ Mpc), allowing us to resolve filamentary and sheet-like structures without introducing excessive fragmentation or noise.
+At this scale, transitional environments such as filaments are sufficiently preserved, enabling a meaningful analysis of gradual quenching and mixed galaxy populations.
+In contrast, smaller smoothing scales (e.g., 2 Mpc) introduce finer detail but can amplify shot noise, while larger scales tend to over-smooth dense structures, diluting cluster cores into their surroundings.
+Moreover, a 4 Mpc filter aligns with the typical width of filaments observed in both simulations and large-scale galaxy surveys, ensuring that our results remain physically interpretable and directly comparable to previous studies.
+Overall, this smoothing length provides an optimal trade-off for capturing the complexity of the cosmic web in a statistically robust and physically meaningful way.